### v1

# Monotonic Path-Specific Effects: Application to Estimating Educational Returns

Abstract
Conventional research on educational effects typically either employs a “years of schooling” measure of education, or dichotomizes attainment as a point-in-time treatment. Yet, such a conceptualization of education is misaligned with the sequential process by which individuals make educational transitions. In this paper, I propose a causal mediation framework for the study of educational effects on outcomes such as earnings. The framework considers the effect of a given educational transition as operating indirectly, via progression through subsequent transitions, as well as directly, net of these transitions. I demonstrate that the average treatment effect (ATE) of education can be additively decomposed into mutually exclusive components that capture these direct and indirect effects. The decomposition has several special properties which distinguish it from conventional mediation decompositions of the ATE, properties which facilitate less restrictive identification assumptions as well as identification of all causal paths in the decomposition. An analysis of the returns to high school completion in the NLSY97 cohort suggests that the payoff to a high school degree stems overwhelmingly from its direct labor market returns. Mediation via college attendance, completion and graduate school attendance is small because of individuals’ low counterfactual progression rates through these subsequent transitions.
Keywords: causal inference, mediation, sequential ignorability, education
1 Introduction
One of the most resilient social scientific findings across a range of national contexts is the strong association between educational attainment and a variety of life outcomes, including earnings, health, social capital, and family stability (Hout, 2012; Chetty et al., 2023). Conventionally, researchers have taken one of two approaches to evaluating the social and economic returns to education: the first employs a “years of schooling” measure of educational attainment (Angrist and Krueger, 1991, 1992; Kane and Rouse, 1993; Card, 1994; Ashenfelter and Zimmerman, 1997; Card, 1999, 2001; Angrist and Chen, 2011), while the second dichotomizes attainment as a point-in-time treatment. This latter approach has been particularly influential in the study of the impact of postsecondary attainment on earnings, where the treatment considered is often an indicator for whether an individual has attended, or graduated from, college (Brand and Xie, 2010; Carneiro et al., 2011; Zimmerman, 2014; Goodman et al., 2017; Smith et al., 2020; Bleemer, 2022; Mountjoy, 2022).
Despite the important insights this literature has made into establishing the causal effect of educational attainment on important social and economic outcomes, extant work has been inattentive to the sequential process by which people make educational transitions (Mare, 1980). [^1] I use the term “educational transition” to refer both to vertical transitions (e.g. enrollment at a secondary or tertiary institution), as well as to the attainment of a qualification at a given level (e.g. high school graduation or BA completion). At the end of high school, individuals decide whether or not to enroll in college. Among college enrollees, only 60% receive a BA within six years of initial college entry (Snyder et al., 2016), with an even lower proportion for low-income students and students of color (Eller and DiPrete, 2018; Zhou and Pan, 2023). Moreover, amidst higher educational expansion in the US, college graduates must increasingly choose whether to enter the labor market or to enroll in post-graduate education. Increasingly, therefore, educational attainment in the US has become a field of multiple levels with sequential transitions, all of which are independently consequential for individuals’ labor market outcomes, and therefore of independent scientific interest.
The sequential nature of educational transitions implies that a causal mediation framework can be employed to study the causal paths by which education’s “value-added” occurs. Specifically, we can consider the first transition in a sequence of educational levels of interest as a treatment variable, $A$, and subsequent transitions as mediators that “transmit” the effects of the treatment and of prior transitions, $M_{k}$ ($1\leq k\leq K$). For example, if we are interested in the total effect of high school completion on earnings, we may ask to what extent this total effect operates indirectly, through the effects of college attendance and college completion (putative mediators) on earnings, or directly, through alternative causal pathways. The insight that the total causal effect of education can be decomposed into its direct and indirect effects opens up a range of important research and policy-oriented questions. For example, tracing to what extent an early-stage educational intervention boosts outcomes such as earnings via its promotion of subsequent educational attainment (its indirect effects), or via earnings directly, would enable policy-makers to discern what drives the intervention’s value and to hone subsequent policy (e.g. Hurwitz and Howell, 2014; Sullivan et al., 2019; Castleman et al., 2020; Bird et al., 2021; Dynarski et al., 2021; Black et al., 2023; Turner and Gurantz, 2024). Relatedly, if the early intervention’s effects are heterogeneous across demographic groups, assessing the intervention’s direct and indirect effects could guide researchers to aspects of the educational experience that either promote or inhibit upward mobility. Nevertheless, prior empirical approaches are not well-suited to answering these questions: a “years-of-schooling” approach captures the direct effect of each additional year of schooling, while the dichotomous approach conflates the direct and indirect effects. [^2] A further strand of literature, especially prominent in labor economics, explores labor market returns to horizontal aspects of differentiation within a given educational level (e.g. college selectivity, as well as specific colleges) or college types (e.g. Cohodes and Goodman, 2014; Goodman et al., 2017; Mountjoy and Hickman, 2021; Chetty et al., 2023; Eller, 2023). While my proposed framework prioritizes the effects of different levels of education, I discuss in my concluding remarks how the framework could be extended to accommodate multivariate mediators.
In this article, I introduce a causal mediation framework for analyzing the effects of educational transitions. For the setting of $K$ ($\geq 1$) monotonic mediators, I develop a general formula that decomposes the total effect of any level of education into $K+1$ monotonic path-specific effects (MPSEs): a direct effect net of $K$ subsequent educational transitions, reflecting the path $A\to Y$, and $K$ mutually exclusive “continuation” or gross effects, reflecting the paths $A\to M_{1}\to Y$, $A\to M_{1}\to M_{2}\to Y$, and $A\rightarrow M_{1}\dots\rightarrow M_{K}\rightarrow Y$. Most importantly, this decomposition exploits a unique characteristic of this empirical setting, in which mediators are characterized by “monotonicity”: that is, where an individual’s potential $k+1$ mediator value is deterministically zero if that individual’s $k$th mediator value is 0. The resultant decomposition of the ATE into $K+1$ monotonic path-specific effects (MPSEs) can be non-parametrically identified under the assumption of sequential ignorability, which allows for the effect of each educational level to be confounded by a distinct set of (observed) intermediate covariates. I introduce several estimation strategies for my proposed decomposition, including a simple linear model-based regression-with-residuals (RWR) procedure, and a non-parametric estimation strategy based on the efficient influence functions (EIFs) of the target parameters (see Chernozhukov et al., 2017; Kennedy, 2022).
This study makes three main contributions. Within the realm of education research, I draw on important work by Heckman et al. (2018), who present a similar decomposition of the effect of schooling over the early life course, but differs in two important respects. First, I provide nonparametric definitions, identification results, and estimation strategies for decomposing the total effect of schooling through its direct and indirect components. Second, my decomposition accommodates the presence of a distinct set of observed intermediate confounders for each transition. While one limitation of my approach is that I assume away the presence of unobserved confounders for each transition, I propose a sensitivity analysis that assesses the robustness of the results to unobserved confounding, under a set of simplifying assumptions.
More broadly, my framework speaks to the burgeoning field of causal mediation analysis in the social, economic, and health sciences, targeted at assessing the causal pathways by which a treatment affects an outcome. While prior literature overwhelmingly focuses on single-mediator decompositions of the ATE, a growing body of work examines mediation estimands in settings with multiple mediators (Avin et al., 2005; Albert and Nelson, 2011; VanderWeele and Vansteelandt, 2014; Lin and VanderWeele, 2017; Miles et al., 2017; Steen et al., 2017; Vansteelandt and Daniel, 2017; Miles et al., 2020). In particular, in the case of two causally ordered mediators, Daniel et al. (2015) show that the ATE can be decomposed into multiple path-specific effects (PSEs), and outline the assumptions under which some of these effects are identified. Most recently, Zhou (2022b) generalized this framework to the case of $K$ mediators, establishing a set of identifiable PSEs and introducing several regression-based, weighting, and semiparametric efficient estimators. I extend this literature by examining a special empirical setting where the mediators are monotonic. Compared with traditional mediation-based decompositions, monotonicity faciliates PSE identification under weaker identification assumptions, enables identification of all of the causal paths in question, as opposed to just a strict subset of them, and further permits a finer-grained decomposition. The general decomposition also extends previous literature on mediation under monotonicity which has focused exclusively on the case of a single mediator (e.g. Zhou, 2022a).
Finally, I also contribute to a growing parallel literature that proposes a range of nonparametric, and semi-parametric efficient estimators for alternative mediation estimands, based on the efficient influence functions (EIFs) of the causal quantities of interest (e.g. Miles et al., 2020; Farbmacher et al., 2022; Zhou, 2022b), as well as to closely-related work that proposes semi-parametric efficient estimators for dynamic treatment effects (Lewis and Syrgkanis, 2020; Viviano and Bradic, 2021; Bodory et al., 2022).
In the following sections, I first introduce the decomposition for the case of a single intermediate educational transition, before discussing the general case of $K$ intermediate transitions and its identification under the assumption of sequential ignorability (Section 2). In Section 3, I introduce a semiparametric estimation strategy for estimating the proposed decomposition, and in Section 4, I illustrate the proposed framework and methods using data from the National Longitudinal Survey of Youth (NLSY97) cohort. Section 5 concludes.
2 Monotonic Path-Specific Effects
2.1 A Single Intermediate Transition
[…]
Using potential outcomes notation, let $M(a)$ denote an individual’s potential value of the mediator if their treatment status were set to $a$, and let $Y(a,m)$ denote that individual’s potential outcome if their treatment and mediator statuses were set to $a$ and $m$, respectively. I assume, as I formalize in the following section, that sequential transitions are characterized by monotonicity; here, this means that individuals who do not complete high school cannot attend college, or $M(0)=0$. This sequential nature of educational transitions therefore implies the following set potential outcomes: $\{Y(1),Y(0),Y(1,0),Y(1,1)\}$. Further, since by the composition assumption $Y(a)=Y(a,M(a))$ (VanderWeele and Vansteelandt, 2009), under monotonicity we have that Y(0)=Y(0,M(0))Y(0)=Y(0,M(0))=Y(0,0)=Y(0,0), and that
$Y(1)$ $=Y(1,0)+M(1)[Y(1,1)-Y(1,0)].$
Thus, the individual total effect of $A$ on $Y$ can be decomposed as
$Y(1)-Y(0)$ $=Y(1,0)-Y(0,0)+M(1)[Y(1,1)-Y(1,0)].$ (1)
Since these individual-level quantities are unidentified, I focus on their population-level analogs. Taking the expectation of Equation 1, we obtain the following decomposition of the ATE (see Zhou, 2022a):
ATE $=\mathbb{E}[Y(1)-Y(0)]$
$=\mathbb{E}[Y(1,0)-Y(0,0)]+\mathbb{E}[M(1)]\mathbb{E}[Y(1,1)-Y(1,0)]+\textup{cov}[M(1),Y(1,1)-Y(1,0)]$ (2)
$=\underbrace{\Delta_{0}}_{A\to Y}+\underbrace{\pi_{\textup{1}}\Delta_{\textup{1}}+\eta_{\textup{1}}}_{A\to M\to Y}.$ (3)
Here, $\Delta_{\textup{0}}$ and $\Delta_{1}$ denote the direct effects of the first and intermediate transitions on the outcome, $A\rightarrow Y$ and $M\rightarrow Y$, respectively, $\pi_{\textup{1}}$ denotes the total effect of the first transition on the intermediate transition $A\to M$, and $\eta_{\textup{1}}$ denotes the covariance between the effect of the initial transition on completion of the second and the effect of the second transition on $Y$. Specifically, $\eta_{\textup{1}}$ is positive if those who would attend college given high school completion (i.e., $M(1)=1$) benefit more from college attendance in terms of their later earnings (i.e., have a larger Y(1,1)Y(1,0)Y(1,1)-Y(1,0)) than those who do not (i.e., $M(1)=0$), and negative if the opposite is true. Meanwhile, the composite term $(\pi_{\textup{1}}\Delta_{1}+\eta_{\textup{1}})$ captures the average indirect effect of the treatment via the intermediate transition ($A\rightarrow M\rightarrow Y$), comprising the sum of (i) the probability of college enrollment if an individual graduated high school, multiplied by the direct of college enrollment, and (ii) the covariance between college enrollment and its direct effect on earnings.
2.2 Generalization to $K$ Intermediate Transitions
I now generalize the approach introduced in the preceding section to the case of $K$ intermediate transitions. As previously, I denote the treatment (“initial transition”) of high school graduation by $A$, and use $M_{1},\dots M_{K}$ to refer to the $K$ subsequent transitions of interest (“intermediate transitions”), where I assume that all of $M_{1},\dots,M_{K}$ are binary and that for any $i<j$, $M_{i}$ temporally precedes $M_{j}$. For instance, we may wish to decompose the total effect of high school completion on earnings via college attendance ($M_{1}$), college completion ($M_{2}$) and graduate school attendance ($M_{3}$). Let an overbar denote a vector of variables, such that $\overline{M}_{k}=(M_{1},M_{2},\dots M_{k})$ and $\overline{1}_{k}=(A=1,M_{1}=1,\dots,M_{k-1}=1)$. Further, let $[K]$ denote the set $\{0,1,\dots,K\}$. In addition, I denote by $X$ a vector of pretreatment confounders of the effect of $(A,\overline{M}_{k})$ on ($M_{k+1},Y)$, and by $\overline{Z}_{k}=(Z_{1},\dots Z_{k})$ a vector of intermediate confounders that may confound the causal effect of $M_{k}$ on ($M_{k+1},Y)$. Using potential outcomes notation, $Y(\overline{1}_{k},m_{k})$ thus denotes an individual’s potential earnings if they completed, possibly contrary to fact, the treatment in addition to k1k-1 intermediate transitions, and then either completed ($m_{k}=1)$ or did not complete ($m_{k}=0$) the $k$th intermediate transition. Similarly, $M_{k+1}(\overline{1}_{k+1})$ denotes an individual’s potential value of the $k+1$th intermediate transition were that individual to complete the treatment as well as $k$ prior intermediate transitions. As is standard in the mediation literature, I make the following composition assumption (VanderWeele and Vansteelandt, 2009):
Assumption 1.
Composition: $Y(\overline{1}_{k},m_{k})=Y(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k})),\forall k\in[K-1],M_{0}\equiv A$.
In words, Assumption 1 states that a person’s potential outcome under $(\overline{1}_{k},m_{k})$ is equal to their potential outcome under $A=1,\dots,M_{k-1}=1,m_{k}$ and under the value $M_{k+1}$ would naturally take under $A=1,\dots,M_{k-1}=1,m_{k}$. I also invoke the following constraint on units’ potential transition values:
Assumption 2.
Monotonicity: $M_{k+1}(M_{k}=0)=0\forall k\in[K-1],M_{0}\equiv A.$
Informally, Assumption 2 (monotonicity) states that an individual’s potential $k+1$th transition value is deterministically 0 if that individual fails to complete the prior ($k$th) transition. It is analogous to a one-sided non-compliance assumption within an instrumental variables (IV) framework, which precludes the presence of both “defiers” as well as “always-takers” principal strata. We can then use this assumption to decompose the ATE of $A$ on $Y$, which I denote by $\tau_{0}$. Specifically, let $\tau_{k}$ denote the gross effect of the $k$th mediator on $Y$, i.e.,
$\tau_{k}=\mathbb{E}[Y(\overline{1}_{k+1})-Y(\overline{1}_{k},0)],$
let $\Delta_{0}$ denote the direct effect of $A$ on $Y$, and let $\Delta_{k}$ denote the direct effect of the $k$th mediator on $Y$, i.e.,
$\Delta_{k}=\mathbb{E}[Y(\overline{1}_{k+1},0)-Y(\overline{1}_{k},0)].$
To explicate my approach, note that the gross effect of the $k$th mediator, $\tau_{k}$, includes not only the direct effect $M_{k}\rightarrow Y$, net of subsequent educational transitions $\Delta_{k}$, but also the indirect effects of $M_{k}$ via subsequent transitions ($M\rightsquigarrow Y$, where a squiggly arrow denotes a combination of multiple paths). This insight motivates us to further decompose $\tau$ into its direct and indirect components. Under the composition assumption, $\tau_{k}$ can be decomposed as
$\tau_{k}=\Delta_{k}+\pi_{k+1}\tau_{k+1}+\eta_{k+1},$ (4)
where
$\pi_{k+1}$ $=\mathbb{E}[M_{k+1}(\overline{1}_{k+1})],$
$\eta_{k+1}$ $=\textup{cov}[M_{k+1}(\overline{1}_{k+1}),Y(\overline{1}_{k+2})-Y(\overline{1}_{k+1},0)].$
For $k=1,\dots,K-1$, iteratively substituting equation 4 into the corresponding expression for $\tau_{k-1}$ yields
$\tau_{0}=\underbrace{\Delta_{0}}_{A\to Y}+\sum_{k=1}^{K}\underbrace{(\Pi_{j=1}^{k}\pi_{j})\Delta_{k}+(\Pi_{j=1}^{k-1}\pi_{j})\eta_{k}}_{\theta_{k}\triangleq A\to M_{1}\ldots\to M_{k}\to Y},$ (5)
where $\Delta_{K}=\tau_{K}$, i.e. $\Delta_{K}$ is a gross or continuation effect, since this latter path is a composite one that contains all residual paths omitted in the decomposition (i.e., through educational transitions subsequent to $K$, if they exist). Thus, the $\theta_{k}$ terms capture how much of the total effect of high school completion flows through each intermediate transition considered (i.e., via college attendance, via college completion, and via graduate school attendance), while $\Delta_{0}$ captures that portion of the total effect that operates directly, net of the $K$ intermediate transitions considered.
2.3 Identification
To identify the causal effects of interest, I rely on a series of sequential ignorability assumptions. While most closely associated with the dynamic treatment effects literature, which rely on observing a complete set of time-varying confounders in order to identify longitudinal effects (see e.g. Lewis and Syrgkanis, 2020; Viviano and Bradic, 2021; Bodory et al., 2022), these assumptions can be transferred to a mediation context, given the fact that the mediators of interest are all causally ordered. As will be discussed in the following section, sequential ignorability identification assumptions are distinct from - and in fact weaker than - the assumptions typically employed in studies of causal mediation.
Before proceeding, I introduce the following shorthands. Let $M_{0}\triangleq A$ and $M_{k}=\emptyset\forall k<0$. In order to estimate the decomposition shown in Equation 5, it suffices to identify the expectation of two types of composite counterfactuals ($Y(\overline{1}_{k},m_{k+1})$ and $M_{k+1}(\overline{1}_{k+1})$), as well as covariance terms of the form $\mathrm{cov}[M_{k+1}(\overline{1}_{k+1}),Y(\overline{1}_{k+2})-Y(\overline{1}_{k+1},0)]$ $\forall k\in[K-1]$. I invoke the following three assumptions:
Assumption 3.
Consistency: for any unit, if $A=a$, $Y=Y(a)$; if $(A,\overline{M}_{k})=\{\overline{1}_{k},m_{k}\}$, then $Y=Y(\overline{1}_{k},m_{k})$ $\forall k\in[K]$, and if $(A,\overline{M}_{k})=\overline{1}_{k+1}$, then $M_{k+1}=M_{k+1}(\overline{1}_{k+1})\forall m_{k+1}\in\{0,1\},\forall k\in[K-1]$.
Assumption 4.
Sequential ignorability: $(M_{1}(1),Y(a))\perp\!\!\!\perp A|X$; $Y(\overline{1}_{k},m_{k})\perp\!\!\!\perp\overline{M}_{k}|X,\overline{Z}_{k},\overline{M}_{k-1}$ and $M_{k+1}(\overline{1}_{k+1})\perp\!\!\!\perp\overline{M}_{k}|X,\overline{Z}_{k},\overline{M}_{k-1},\forall m_{k}\in\{0,1\},\forall k\in\{1,\dots,K\},M_{0}\equiv A$.
Assumption 5.
Positivity: $p_{A|X}(a|x)>\epsilon>0$, $p_{M_{k}|X,A,\overline{Z}_{k},\overline{M}_{k-1}}(m_{k}|x,a,\overline{z}_{k},\overline{m}_{k-1})>\epsilon>0$ $\forall k\in[K]$.
Assumption 3 (consistency) states that a unit’s observed outcome equals its potential outcome under a given treatment sequence. Note that under under the Assumption 1 (Composition), if $Y=Y(\overline{1}_{k},m_{k})$, then $Y=Y(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k}))=Y(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k})),$ $\dots,M_{K}(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k}),\dots M_{K-1}((\overline{1}_{k},m_{k}),M_{k+1}(\overline{1}_{k},m_{k}),\dots,M_{k-2}(\dots))$.
In plain words, the KkK-k mediators after mediator $k$ all take their natural levels. Assumption 4 (sequential ignorability) is the no unmeasured confounding assumption for the treatment and all mediators. It is considered plausible when sufficient pre-treatment and intermediate covariates $(X,\overline{Z}_{K})$ are collected. Finally, Assumption 5 (positivity) requires that treatment and mediator assignment is not deterministic. Under Assumptions 3-5, $\mathbb{E}[Y(\overline{1}_{k},m_{k})]$ and $\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]$ are identified, respectively, as
$\mathbb{E}[Y(\overline{1}_{k},m_{k})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k}]\big{[}\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x)$ (6)
$\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[M_{k+1}|x,\overline{z}_{K},\overline{1}_{k+1}]\big{[}\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x)$ (7)
For a proof of the above formulas, see Robins (1986). The covariance ($\eta_{k}$) components in the decomposition are then identified as the “residual” terms such as in Equation 4, which follows directly from the fact that all other components in these equations are identified. Thus, for $k\in\{1,\dots K\}$, we can identify $\eta_{k}$ as
$\eta_{k}=\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k}.$
2.4 A Comparison with Conventional Mediation Analysis with Multiple Causally Ordered Mediators
The above decomposition has an analog in the context of a mediation–based decomposition of the ATE with multiple ordered mediators, but differs from conventional mediation analysis in important ways. To illustrate the differences, consider a binary treatment, $A$, an outcome of interest, $Y$, and a vector of pretreatment covariates, $X$, and let $M_{1},M_{2},\ldots M_{K}$ denote $K$ causally ordered mediators, assuming that for any $i<j,M_{i}$ precedes $M_{j}$, as above. Moreover, let an overbar denote a vector of variables, so that $\bar{M}_{k}=(M_{1},M_{2},\ldots M_{k})$$\bar{m}_{k}=(m_{1},m_{2},\ldots m_{k})$, and $\bar{a}_{k}=(a_{1},a_{2},\ldots a_{k})$. Using the potential outcomes notation as above, we can define the following expectation of a nested counterfactual,
$\psi_{a\bar{a}_{k}}\triangleq\mathbb{E}\left[Y(a,\bar{M}_{k}(\bar{a}_{k}))\right],$
where $\bar{M}_{k}(\bar{a}_{k})\triangleq\big{(}\bar{M}_{k-1}(\bar{a}_{k-1}),M_{k}(a_{k},\bar{M}_{k-1}(\bar{a}_{k-1}))\big{)},\forall k\in[K]$. Under Pearl’s (2009) nonparametric structural equation model (NPSEM), Zhou (2022b) demonstrates that the ATE of $A$ on $Y$ can be decomposed into $K+1$ identifiable path-specific effects (PSEs) corresponding to each of the causal paths $A\rightarrow Y$ and $A\rightarrow M_{k}\leadsto Y$ $(k\in[K])$:
$\mathrm{ATE}=\psi_{\overline{1}}-\psi_{\overline{0}}=\underbrace{\psi_{1,\overline{0}_{K},}-\psi_{\overline{0}_{K+1}}}_{A\rightarrow Y}+\sum_{k=1}^{K}\underbrace{\big{(}\psi_{\overline{1}_{k+1},\underline{0}_{k+1}}-\psi_{\overline{1}_{k},\underline{0}_{k+1}}}_{A\rightarrow M_{k}\rightsquigarrow Y}\big{)}.$ (8)
This decomposition holds algebraically when Assumption 2 does not hold (i.e., when the mediators are not monotonic). In contrast, the monotonic characteristic of the proposed decomposition leads to several important differences. First, the PSE decomposition of the ATE in general mediation settings is not algebraically unique, and thus the PSEs defined under alternative decompositions will differ if the effects of the treatment and each mediator vary across levels of the other mediators. In fact, depending on the order in which the paths $A\rightarrow Y$ and $A\rightarrow M_{k}\rightsquigarrow Y$ are considered, there are $(K+1)!$ identifiable different ways of decomposing the ATE; the decomposition shown in Equation 8 is just one such decomposition. Consider the case of two causally dependent mediators. In this setting, the causal pathway $A\to M_{2}\rightsquigarrow Y$ can be defined with respect to four different combinations of levels of the treatment and first mediator: under (i) $a=1$ and $M_{1}(1)$, (ii) $a=1$ and $M_{1}(0)$, (iii) $a=0$ and $M_{1}(1)$, or (iv) $a=0$ and $M_{1}(0)$. By contrast, as a direct consequence of monotonicity, the MPSE decomposition is the unique PSE decomposition of the ATE.
Second, for general PSE decompositions of the ATE, the set of identifiable decompositions is merely a small subset of the total number of decompositions that hold algebraically (see Avin et al., 2005). In particular, the identifiable decomposition does not enable us to disentangle the mediating effects of $M_{k}$ that are direct (net of subsequent mediators) and indirect (through different combinations of subsequent mediators). For example, in the case of two causally dependent mediators, to assess the mediating role of $M_{1}$, only the composite path $A\rightarrow M_{1}\rightsquigarrow Y=(A\rightarrow M_{1}\rightarrow Y)+(A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y)$ is identified. By contrast, mediator monotonicity permits a finer-grained decomposition of the ATE: each PSE is identified. In the case of two causally dependent mediators, for example, the causal path $A\rightarrow M_{2}\rightarrow Y$ is zero, and as a result, each of the paths $A\to Y$, $A\rightarrow M_{1}\rightarrow Y$ and $A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y$ are identifiable. Figure 1 illustrates the causal pathways defined and identified under the proposed decomposition in the case of two monotonic mediators.
Finally, the sequential ignorability assumption required to identify the MPSE decomposition is weaker than those required to identify a generic PSE decomposition of the ATE. Specifically, the latter requires Pearl’s (2009) non-parametric structural equation model (NPSEM), which stipulates that $\left(M_{k+1}\left(a_{k+1},\bar{m}_{k}\right),\ldots M_{K}\left(a_{K},\bar{m}_{K-1}\right),Y\left(a_{K+1},\bar{m}_{K}\right)\right)\perp\!\!\!\perp M_{k}\left(a_{k},\bar{m}_{k-1}^{*}\right)\mid X,A,\bar{M}_{k-1},\forall k\in[K].$ This assumption, sometimes referred to as the “cross-world” independence assumption, is stronger than the sequential ignorability assumption (4) required to identify the MPSE decomposition since it rules out the existence of confounders of the mediators, be they observed or unobserved. By contrast, the MPSE decomposition identification results accommodate observed intermediate confounding without altering the substance of the decomposition.
Figure 1: Causal Relationships with Two Monotonic Mediators Shown in a Directed Acyclic Graph (DAG) and the 3 Monotonic Path Specific Effects (MPSEs). AA denotes an initial transition of interest, $Y$, an outcome, and $M_{1}$ and $M_{2}$ are two causally ordered, monotonic mediators. The set $(X,Z_{1},Z_{2})$ captures pre-treatment and intermediate confounders.
3 Semiparametric, EIF-Based Estimation
The identification results outlined above suggest that the proposed decomposition can be estimated via several approaches, including outcome-based modeling, models for the treatment and mediators via inverse probability weighting, as well as doubly robust approaches. Parametric procedures are attractive because of their conceptual simplicity and ease of implementation: in Supplementary Material A I show how, under a set of linear models for the outcome and mediators, the $\theta_{k}$ components in Equation 5 can be read off from simple functions of coefficients in these linear models. However when $X$ and $\overline{Z}_{K}$ are high-dimensional, parametric estimators which require a user-defined specification of the data-generating process may suffer biases resulting from model misspecification. In order to reduce model dependency, in this section I provide a nonparametric estimation approach which draws on a debiased machine learning (DML) approach. My DML approach is characterized by two components: first, the use of a Neyman orthogonal estimating equation based on the efficient influence function (EIF) for the target parameters, which makes estimates of the parameter “locally robust” to estimates of the nuisance functions; second, the use of a $K$-fold cross-fitting algorithm (Chernozhukov et al., 2017).
Let $O=(X,A,\overline{Z}_{K},\overline{M}_{K},Y)$ denote the observed data, and $\mathcal{P}$ a nonparametric model over $O$ wherein all laws satisfy the positivity assumption described in Section 2. Before proceeding, I define the following auxiliary functions, as introduced in Section 2: $\psi_{km_{k}}\triangleq\mathbb{E}[Y(\overline{1}_{k},m_{k})]$ and $\phi_{k}\triangleq\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]$, for all $k\in[K],M_{0}\triangleq A$. Using the identification results given in Section 2, $\psi_{km_{k}}$ can be written in terms of expectations of observed data:
$\psi_{km_{k}}=\mathbb{E}_{X}\mathbb{E}_{Z_{1}|X,1}\ldots\mathbb{E}_{Z_{k}|X,\overline{Z}_{k-1},\overline{1}_{k}}\mathbb{E}[Y|X,\overline{Z}_{k},\overline{1}_{k},m_{k}].$
For each $j\in[k]$, we can thus define $\mu_{jm_{k}}^{k}\left(X,\bar{Z}_{k}\right)$ iteratively as
$\mu_{km_{k}}^{k}\left(X,\bar{Z}_{k}\right)$ $\triangleq\mathbb{E}\left[Y\mid X,\bar{Z}_{k},\overline{1}_{k},m_{k}\right],$
$\mu_{jm_{k}}^{k}\left(X,\bar{Z}_{j}\right)$ $\triangleq\mathbb{E}\left[\mu_{j+1m_{k}}^{k}\left(X,\bar{Z}_{j+1}\right)\mid X,\bar{Z}_{j},\overline{1}_{j+1}\right]\forall j\in[k-1].$
Further, let $\pi_{km_{k}}\big{(}X,\overline{Z}_{k}\big{)}\triangleq\mathrm{Pr}[M_{k}=m_{k}\mid X,\overline{Z}_{k},\overline{1}_{k}]\forall k\in[K]$, and $\pi_{01}\big{(}X\big{)}\triangleq\text{Pr}[A=1|X]$. The efficient influence function (EIF) of $\psi_{km_{k}}$ is closely related to the EIF for the g-formula, and can be written as
$\psi_{km_{k}}(O)=\sum_{j=0}^{k+1}\varphi_{j}(O),$
where
$\varphi_{0}(O)$ $=\mu_{0m_{k}}^{k}(X)-\psi_{km_{k}}$
$\varphi_{j}(O)$ $=\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\prod_{l=1}^{j-1}\frac{M_{l}}{\pi_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\left(\mu_{jm_{k}}^{k}\left(X,\overline{Z}_{j}\right)-\mu_{j-1m_{k}}^{k}\left(X,\bar{Z}_{j-1}\right)\right),\quad j\in\{1\dots,k\}$
$\varphi_{k+1}(O)$ $=\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\frac{\mathbb{I}(M_{k}=m_{k})}{\pi_{km_{k}}\big{(}X,\overline{Z}_{k}\big{)}}\prod_{k=1}^{k-1}\frac{M_{l}}{\pi_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\left(Y-\mu_{km_{k}}^{k}\left(X,\bar{Z}_{k}\right)\right).$
[…]
$\hat{\psi}_{km_{k}}^{\text{eif}}$ $=\mathbb{P}_{n}\bigg{[}\frac{A}{\hat{\pi}_{01}\big{(}X\big{)}}\left(\frac{\mathbb{I}(M_{k}=m_{k})}{\hat{\pi}_{km_{k}}\big{(}X,\overline{Z}_{k}\big{)}}\prod_{k=1}^{k-1}\frac{M_{l}}{\hat{\pi}_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\left(Y-\hat{\mu}_{km_{k}}^{k}\left(X,\bar{Z}_{k}\right)\right)$ (9)
$+\sum_{j=1}^{k}\frac{A}{\hat{\pi}_{01}\big{(}X\big{)}}\left(\prod_{l=1}^{j-1}\frac{M_{l}}{\hat{\pi}_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\left(\hat{\mu}_{jm_{k}}^{k}\left(X,\overline{Z}_{j}\right)-\hat{\mu}_{j-1m_{k}}^{k}\left(X,\bar{Z}_{j-1}\right)\right)$
$+\hat{\mu}_{0m_{k}}^{k}(X)\bigg{]}.$
A similar EIF-based estimator can be used for $\phi_{k}$ to estimate the $\pi_{k}$ terms in Equation 5. This estimator is based on the following nuisance functions for estimation (see Supplementary Material I for details):
$\gamma_{k}\left(X,\bar{Z}_{k}\right)$ $\triangleq\mathbb{E}\left[M_{k+1}\mid X,\bar{Z}_{k},\overline{1}_{k+1}\right],$
$\gamma_{j}\left(X,\bar{Z}_{j}\right)$ $\triangleq\mathbb{E}\left[\gamma_{j+1}\left(X,\bar{Z}_{j+1}\right)\mid X,\bar{Z}_{k},\overline{1}_{j+1}\right]\forall j\in[k-1].$
Next, following Kennedy (2022, p. 15), let $\mathbb{IF}:\Psi\rightarrow L_{2}(\mathbb{P})$ denote the operator mapping the functionals $\{\Delta_{k},\pi_{k},\eta_{k}\}:\mathcal{P}\rightarrow\mathbb{R},$ $\forall\in[K]$ to their respective influence functions under the nonparametric model $\mathcal{P}$. Because the $(\Delta_{k},\tau_{k}$) components of the decomposition are linear in $\psi_{km_{k}}$, by linearity of the EIF, $(\mathbb{IF}(\Delta_{k}),\mathbb{IF}(\tau_{k})$) can be expressed as linear combinations of $\varphi_{km_{k}}(O)$. In particular, $\mathbb{IF}(\tau_{k})=\varphi_{(k+1)1}(O)-\varphi_{k,0}(O)$ and $\mathbb{IF}(\Delta_{k})=\varphi_{(k+1)0}(O)-\varphi_{k,0}(O)$. The EIFs of $\eta_{k}$ and $\theta_{k}$, $\forall k\in[K]$ under $\mathcal{P}$ are derived as in Proposition 1:
Proposition 1.
The EIFs of $\eta_{k}$, $\theta_{k}$ $\forall k\in[1,\dots,K]$ under $P$ are given, respectively, by
$\mathbb{IF}(\eta_{k})$ $=\mathbb{IF}(\tau_{k-1})-\mathbb{IF}(\Delta_{k-1})-\tau_{k}\mathbb{IF}(\pi_{k})-\pi_{k}\mathbb{IF}(\tau_{k}),$
$\mathbb{IF}(\theta_{k})$ $=\mathbb{IF}(\Delta_{k})\prod_{j=1}^{k}\pi_{j}+\Delta_{k}\sum_{j=1}^{k}\mathbb{IF}(\pi_{j})\prod_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{k}\pi_{l}+\mathbb{IF}(\eta_{k})\prod_{j=1}^{k-1}\pi_{j}+\eta_{k}\sum_{j=1}^{k-1}\mathbb{IF}(\pi_{j})\prod_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{k-1}\pi_{l},$
for $k\in\{1,\dots K\}$, with $\theta_{0}=\Delta_{0}$, and where $\mathbb{RIF}(\phi)=\mathbb{IF}(\phi)+\phi,$ denotes the recentered EIF of a parameter (about the truth). Their corresponding EIF-based estimators are (see Supplementary Material I for derivations):
$\hat{\eta}_{k}^{\text{eif}}$ $=\mathbb{\widehat{\mathbb{RIF}}}(\tau_{k-1})-\widehat{\mathbb{RIF}}(\Delta_{k-1})-\hat{\tau}_{k}\widehat{\mathbb{RIF}}(\pi_{k})-\hat{\pi}_{k}\widehat{\mathbb{RIF}}(\tau_{k})+\hat{\pi}_{k}\hat{\tau}_{k},$
$\hat{\theta}_{k}^{\text{eif}}$ $=\widehat{\mathbb{RIF}}(\Delta_{k})\prod_{j=1}^{k}\hat{\pi}_{j}+\hat{\Delta}_{k}\sum_{j=1}^{k}\widehat{\mathbb{RIF}}(\pi_{j})\prod_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{k}\hat{\pi}_{l}+\widehat{\mathbb{RIF}}(\eta_{k})\prod_{j=1}^{k-1}\hat{\pi}_{j}+\hat{\eta}_{k}\widehat{\mathbb{RIF}}(\pi_{j})\prod_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{k-1}\hat{\pi}_{l}$
$\;-k\hat{\Delta}_{k}\prod_{j=1}^{k}\hat{\pi}_{j}-(k-1)\hat{\eta}_{k}\prod_{j=1}^{k-1}\hat{\pi}_{j}.$
where $\widehat{\mathbb{RIF}}(\phi)=\widehat{\mathbb{IF}}(\phi)+\phi,$ and $\widehat{\mathbb{IF}}(\phi)$ denotes the influence function of a parameter evaluated at estimates of its component nuisance functions (see Supplementary Material I for derivations).
When machine learning estimators are used to compute the nuisance functions, in order to ensure the convergence rates outlined in Proposition 2 below, one could assume Donsker-type conditions for the nuisance function estimators, which restricts the set of estimators available to use. Alternatively, to expand the class of estimators that can be used for estimating the nuisance functions, sample-splitting can be used. In particular, Chernozhukov et al. (2017) suggest a “cross-fitting” procedure, which comprises the following steps: (1) Randomly split data into $J$ folds: $\{S_{1},...S_{J}\}$; (2) For each fold $S_{j}$, use the remaining (j1j-1) folds (training sample) to fit a flexible machine-learning model for each of the nuisance functions involved in the estimating equations; (3) For each observation in $j$ (estimation sample), use estimates of the above models to construct a set of estimated RIF functions for $\Delta_{k}\forall k\in\{0,\dots,K-1\}$, and for $(\pi_{k},\tau_{k},\eta_{k},\theta_{k})\forall k\in[K]$; (4) Compute an estimate of the decomposition components by averaging the estimated RIF functions across all subsamples $S_{1}$ through $S_{J}$. When all nuisance functions are estimated via data-adaptive methods and cross-fitting, the semiparametric efficiency of $\theta_{k}^{\mathrm{rEIF}}$ is given in the following proposition:
Proposition 2.
Semiparametric efficiency. Under Assumption 5, and under suitable regularity conditions (e.g. Chernozhukov et al., 2018), then $\hat{\theta}_{k}^{\text{eif}}$ is semiparametric efficient if $\sum_{j=k}^{k+1}\bigg{[}\sum_{l=0}^{j}R_{n}\big{(}\hat{\pi}_{l1})R_{n}\big{(}\hat{\mu}_{l0}^{j})\bigg{]}+\sum_{j=0}^{k-1}\bigg{[}R_{n}\big{(}\hat{\pi}_{j1})R_{n}\big{(}\hat{\mu}_{j0}^{k-1})+R_{n}\big{(}\hat{\pi}_{j1})R_{n}\big{(}\hat{\mu}_{j1}^{k-1})\bigg{]}+\sum_{j=0}^{k}\bigg{[}\sum_{l=0}^{j}R_{n}\big{(}\hat{\pi}_{l1})R_{n}\big{(}\hat{\gamma}_{l}^{j})\bigg{]}=o(n^{-1/2})$, where $R_{n}\big{(}\cdot\big{)}$ denotes a mapping from a nuisance function to its $L_{2}(P)$ convergence rate, and where $\hat{\mu}_{l0}^{K+1}\triangleq\hat{\mu}_{l1}^{K}$.
[…]
When data-adaptive methods are used to estimate the nuisance functions, inference on all components can be conducted via the variance of the empirical analog of the EIF, i.e. $\mathbb{P}_{n}[\big{(}\hat{\psi}_{km_{k}}^{\text{EIF}}\big{)}^{2}]/n$. For example, inference on $\tau_{1}$ can be conducted by estimating $\mathbb{P}_{n}[\big{(}\hat{\psi}_{11}^{\text{EIF}}-\hat{\psi}_{10}^{\text{EIF}}\big{)}^{2}]/n$.
4 Empirical Analysis
To illustrate my approach empirically, I draw on data from the National Longitudinal Survey of Youth 1997 (NLSY97). I parse out the direct effect of high school graduation on adult earnings and its indirect or continuation effects via (i) college attendance, (ii) college graduation, and (iii) graduate school attendance. My analytic sample comprises $N=7,305$ respondents.
I construct four types of variables: educational transitions, adult earnings, a set of confounders for the effect of high school graduation on subsequent transitions and earnings, and a single set of intermediate confounders for the effect of college completion on subsequent transitions and earnings. My educational transition variables contain a binary treatment denoting whether a respondent had graduated high school by age 22, and three binary mediators denoting whether the respondent had attended a 4-year college by age 22, whether the respondent had received a BA degree by age 29, and whether the respondent had enrolled in a graduate level program by age 29, respectively. I assume that all individuals who make a given educational transition have made all previous educational transitions. Thus, by construction, my coding strategy disallows for cases which violate the monotonicity assumption. [^3] Assuming away cases in which an individual makes a particular educational transition without having made all previous transitions serves as a reasonable approximation to reality. Among the set of individuals who have non-missing earnings information in the NLSY97 (i.e., those who comprise my analytic sample), 94% of individuals observed to attend graduate school by age 29 also completed a BA by age 29; 93% of respondents who completed a BA by age 29 had attended a 4-year college by age 22 (6% of those who completed a BA by age 29 first attended a 4-year college between ages 23 and 26 inclusive), and 99% of respondents who attended a 4-year college by age 22 had also completed high school. My outcome of interest is logged average annual earnings at ages 32-36, which I define to be the (logged) average of a respondent’s self-reported wage, salary income, and business income. Earnings are adjusted for inflation to 2023 dollars using the personal consumption expenditures (PCE) index. After dropping respondents with missing earnings information, I accommodate those with zero earnings by adding a small constant of \$1,000\\$1,000 to observed earnings (though in Supplementary Material F, I replicate my main analyses under alternative definitions of earnings).
In an effort to satisfy the sequential ignorability assumption (Assumption 4), I include a large array of covariates in my models. This set of covariates is more expansive than those used in previous, observational studies of returns to education (see in particular Scott-Clayton and Wen, 2019). In particular, in addition to including information on respondent demographics (gender, race, ethnicity, age in 1997), and observed pre-college performance such as overall high school GPA and test score on the Armed Services Vocational Aptitude Battery (ASVAB), I include detailed information on socioeconomic background. Since my proposed decomposition also facilitates the inclusion of a distinct set of observed intermediate confounders for each transition, I include two postsecondary characteristics ($Z$) to adjust for confounders of the effect of BA completion and graduate school attendance on earnings: field of study and college GPA. To assess the robustness of my main conclusion to forms of unobserved confounding, in Supplementary Material C, I produce a set of “bias-corrected” estimates of the decomposition components under certain assumptions about the nature of the confounding.
A large proportion (just under $50\%$) of respondents are missing information on covariates $X$ and $Z$. For my main analyses, I impute missing values on these covariates via multiple imputation to increase efficiency, but in Supplementary Material E, I replicate these analyses restricted to the sample of respondents with complete information. This exercise produces substantively similar results (for covariate means for each of these analytic samples, see Supplementary Material D). After constructing the analytical sample, I apply both the DML estimator described in Section 3 as well as a parametric, regression-with-residuals (RWR) algorithm (described in Supplementary Material A) to implement the proposed decomposition. For the DML approach, I estimate all nuisance functions, using a super learner composed of the Lasso and random forest and, following Chernozhukov et al. (2017), use five-fold cross-fitting. All weights involved in computing the rEIFs are censored at their $1st$ and $99th$ percentiles. Supplementary Material H gives further details about the particular models required given my assumed data generation process.
Figure 2 shows my estimates of the average total effect (ATE) on log earnings and its direct and continuation components under both the DML and RWR procedures. Both procedures return similar estimates, though deviate in the estimated magnitude of MPSE $\theta_{1}$, and DML estimates come expectedly with a significantly greater amount of precision. The first column shows that the estimated ATE of graduating high school on log earnings under DML (RWR) is $0.67$ ($0.63$), which implies an earnings premium of approximately $96\%$. The next two columns indicate that the vast majority ($69\%$ under DML and $75\%$ under RWR) of the ATE operates directly, i.e. net of college attendance, BA completion and graduate school attendance (MPSE $\theta_{0}$, $A\rightarrow Y$). Specifically, high school graduates who do not proceed to college can be expected to earn on average $0.46$ ($0.47)$ log earnings more than high school non-completers under DML (RWR), an earnings premium of $59\%$.
While the majority of the ATE is explained by the direct effect, a non-trivial portion occurs through mediation effects through later transitions. Under DML, the continuation effects of high school graduation via college attendance without BA completion (MPSE $\theta_{1}$, $A\rightarrow M_{1}\rightarrow Y$) and via BA completion without graduate school participation (MPSE $\theta_{2}$, $A\rightarrow M_{1}\to M_{2}\rightarrow Y$) both mediate roughly $15\%$ of the ATE, and correspond to an earnings premium of approximately $10\%$. The RWR estimate of $\theta_{1}$ is notably lower at $0.03$ and is also imprecisely estimated. Under both estimation procedures, the continuation effect via graduate school attendance ($A\rightarrow M_{1}\rightarrow M_{2}\rightarrow M_{3}\rightarrow Y$) is very small and fails to reach conventional levels of significance. In sum, the total effect of high school graduation on earnings is determined overwhelmingly by its direct effect on earnings.
Table 1 shows DML and RWR estimates of the various components (the direct effects ($\Delta_{k}$), probabilities ($\pi_{k}$) and covariance terms ($\eta_{k}$)) that constitute the continuation effects $\theta_{k}$. Several points are of note. First, the components in the table offer insights into the economic and educational returns to different educational stages. The direct effects of each educational transition ($\Delta_{k}$) are highly variable: they are largest for high school graduation and for college completion (both at $0.46$ under DML), and lowest for college attendance and graduate school participation (at $0.2$ and $0.12$, respectively, under DML). Note that the payoff to graduate school attendance could be depressed by the fact that I observe individuals at a maximum age of only 36, if graduate school earnings premia materialize only much later in the life course. The counterfactual continuation probabilities ($\pi_{k}$) also provide insight into barriers in educational participation. In particular, even if an individual were to complete high school (possibly contrary to fact), that individual would have under a $50\%$ chance of continuing to a 4-year college without further intervention to increase individuals’ college application, admissions and enrollment rates. Further, even if individuals were to counterfactually both complete high school and attend a 4-year college, only a very small proportion ($\pi_{1}\cdot\pi_{2}=0.24$) would be expected to complete their BA degree without further intervention at the college-level.
Second, the fine-grained nature of the MPSE decomposition enables us to trace the continuation effects to their constituent components. In particular, while the direct effect of high school completion is comparable to the direct effect of BA graduation on earnings, suggesting an earnings premium of $59\%$ relative to college attendance without completion, the continuation effect via BA completion that it informs (MPSE $\theta_{2}$, $A\rightarrow M_{1}\to M_{2}\rightarrow Y$) only mediates a small amount of the overall ATE because because $\theta_{2}$ is approximately (plus the small value of $\eta_{2}$) equal to $\Delta_{2}$ scaled by the product $\pi_{1}\cdot\pi_{2}=0.24$. In words, despite the relatively large direct effect of BA completion on earnings, given individuals’ low counterfactual probability of BA completion, this transition is not an important mediating pathway of the total effect of high school completion on earnings. The result is that college attendance without completion mediates high school graduation’s earnings effects as much as BA completion, despite the fact that college attendance without completion yields a much smaller earnings return for high school graduates than BA completion without graduate school attendance does for college enrollees.
One instructive point of comparison for these results are instrumental variable (IV) estimates of returns to years of schooling, typically estimated in the range of $6\%$ to $12\%$ (Angrist and Krueger, 1991, 1992; Kane and Rouse, 1993; Card, 1994; Ashenfelter and Zimmerman, 1997; Angrist and Chen, 2011). While my estimate of the overall return to high school graduation ($\tau_{0})$ could appear large in this light, several factors could reconcile this difference. First, $\tau_{0}$ captures the direct and continuation effects of high school completion (whereas IV estimates of schooling returns capture schooling’s direct effects). Further, $\tau_{0}$ captures the effect of multiple additional years of schooling (as the high school graduates and high school non-completers that form the comparison group differ by multiple years of schooling), as opposed to a single year’s additional return. In fact, we can more directly compare my DML estimate of the direct return to high school graduation ($\Delta_{0}$) of $0.46$ (corresponding to an earnings premium of $58\%$) using the fact that, in the NLSY97, high school non-completers attained on average $3.7$ fewer years of schooling than high school completers. An IV estimate of $12\%$, for example, would therefore imply an earnings return to $3.7$ additional years of approximately $52\%$ - broadly in line with my result. Still, to assess the robustness of the above findings to potential violations of Assumption 4 (Sequential Ignorability), I implement a sensitivity analysis in Supplementary Material C. Under the stated assumptions about the pattern of unobserved confounding, my primary finding that the ATE of high school graduation is overwhelmingly mediated via its direct effect remains highly robust to unobserved confounding.
Figure 2: Decomposition of the Average Total Effect (ATE) of High School Graduation on Logged Earnings via Debiased Machine-Learning (DML) and Regression-With-Residuals (RWR).
Table 1: Direct Effects ($\Delta_{k}$), Probabilities ($\pi_{k}$) and Covariance Terms ($\eta_{k}$) Involved in Decomposition via Debiased Machine-Learning (DML) and Regression-With-Residuals (RWR).
$\Delta_{0}$ $\Delta_{1}$ $\Delta_{2}$ $\Delta_{3}$ $\pi_{1}$ $\pi_{2}$ $\pi_{3}$ $\eta_{1}$ $\eta_{2}$ $\eta_{3}$
DML 0.462 0.200 0.463 0.122 0.427 0.554 0.315 0.006 0.005 -0.016
(0.059) (0.034) (0.046) (0.029) (0.009) (0.015) (0.022) (0.007) (0.007) (0.009)
RWR 0.469 0.117 0.491 0.160 0.374 0.515 0.219 -0.016 0.071 0.017
(0.115) (0.082) (0.097) (0.113) (0.015) (0.066) (0.018) (0.007) (0.015) (0.046)
{justify}
Note: The $\Delta_{k}$ parameters capture the average effect of completing the $k$th mediator but no subsequent mediator on earnings, relative to completing the k1k-1th mediator. For instance, $\Delta_{0}$ denotes the effect of completing high school ($M_{1}$) but not attending college nor, under Assumption 2, completing any subsequent mediators, relative to attending high school but not completing it ($M_{0}\equiv A$). The $\pi_{k}$ terms capture the average of individuals’ counterfactual completion status of the $k$th mediator under completion of all prior mediators $M_{0},\dots M_{k-1}$. For example, $\pi_{1}$ denotes individuals’ average counterfactual college attendance, after - possibly contrary to fact - their completion of high school. Finally, the $\eta_{k}$ terms refer to the covariance between individuals’ own counterfactual completion status of the $k$th mediator, and their own “gross” effect of completing the $k$th mediator on earnings. To recall, the “gross” effect of the $k$th mediator captures the effect of completing that mediator, relative to completing only the k1k-1th mediator, irrespective of whether that effect operates directly (net of subsequent mediators) or via subsequent transitions. For example, $\eta_{1}$ denotes the covariance between each individual’s counterfactual college attendance status and their gross effect of college attendance on earnings.
5 Conclusion
In this article, I have developed a causal mediation framework for analyzing education effects on earnings. First, I have demonstrated that the total effect of any level of education can be decomposed into a direct effect and $K$ mutually exclusive “continuation” effects. All of these effects are identifiable under the assumption of sequential ignorability. Importantly, this property allows for the effect of each educational transition to be confounded by a distinct set of observed covariates - a property which allows for weaker identification conditions compared with conventional mediation-based decompositions of the ATE (Miles et al., 2017; Zhou, 2022b).
Although my empirical motivation is the estimation of educational returns, the proposed framework applies widely to a range of demographic and organizational settings characterized by “state dependency” between treatment and mediators. This characteristic is particularly salient in demographic phenomena, which often involve sequential transitions over the life course. Certain demographic events are rigid in their monotonicity as a result of their definition. For example, researchers may be interested in discerning the degree to which positive effects of marriage on outcomes such as earnings and life satisfaction are undermined by the negative effects of divorce and separation (and, in turn, their mitigation via re-marriage) (Kenney, 2004; Sweeney and Phillips, 2004). Divorce can be “attained” only by individuals who are already married. Similarly, the effect of parenthood on earnings can be seen as operating directly, through the effect of having a first child net of subsequent children, as well as operating indirectly through the effects of having multiple children, transitions which are clearly monotonic in nature. A similar perspective may be taken in a criminal justice context: the total effect of early-stage police contact (such as being searched for contraband) on educational and socio-psychological outcomes can be decomposed into path-specific effects via subsequent arrest and incarceration (Weaver and Lerman, 2010; Kirk and Sampson, 2013; Sugie and Turney, 2017).
Finally, although in this paper I have considered a decomposition of the average treatment effect for the case of binary monotonic mediators, as shown in Supplementary Material G, the framework could straightforwardly be extended to accommodate categorical transitions. Given the heterogeneity of higher-educational trajectories in the US, such an extension would prove useful for modeling the relative payoffs to distinct educational pathways.
SUPPLEMENTARY MATERIAL
Supplementary Materials (To Appear Online):
This document (PDF) contains: (A) details about a parametric, regression-with-residuals estimation procedure; (B) a simulation study, (C) a sensitivity analysis, (D) further information on sample construction and the data used, (E-F) results with alternative sample restrictions and definitions of earnings; (G) a discussion of extending the decomposition to multivalued, discrete mediators; (H) a description of the EIFs used in the empirical illustration, and (I) proofs and technical details.
Supplemental Materials (to appear online)
1 Parametric, regression-with-residuals (RWR) estimation
In this section, I propose a linear regression-with-residuals (RWR) approach for the MPSE decomposition. The approach relies on two steps. The first involves residualizing pre-treatment confounders with respect to their marginal means, and intermediate confounders on all causally prior confounders, ie., $X^{\perp}\triangleq X-\mathbb{E}[X]$, and $Z_{k}^{\perp}\triangleq M_{k-1}\big{[}Z_{k}-\mathbb{E}[Z_{k}\mid X,\overline{Z}_{k-1},M_{k-1}=1]\big{]}$ for all $k\in[K]$, $M_{0}\triangleq A$. For now, we are agnostic about the functional form used for $\mathbb{E}[Z_{k}\mid X,\overline{Z}_{k-1},M_{k-1}=1]$. The second step involves fitting three sets of models. The first is simply a model for the outcome given pre-treatment covariates and the treatment, namely,
$\mathbb{E}[Y\mid X,A]$ $=\lambda_{0}+\lambda_{1}A+\alpha_{1}^{T}X^{\perp}+\alpha_{2}^{T}AX^{\perp};$ (10)
The second is a set of models for the outcome given covariates, the treatment and $M_{k}$ for all $k\in[K]$, i.e.,
$\mathbb{E}[Y|X,\overline{Z}_{k},A,\overline{M}_{k}]=$ $\beta_{k,0}+c_{k,0}A+\sum_{j=1}^{k}\beta_{k,j}M_{j}+\eta_{k,1}^{\top}X^{\perp}+c_{k,1}AX^{\perp}+\sum_{j=1}^{k-1}\eta_{k,j}^{T}M_{j}X^{\perp}$ (11)
$+\sum_{j=1}^{k}\gamma_{k,j}^{T}Z_{j}^{\perp}+\sum_{j=1}^{k-1}M_{j}\sum_{l=1}^{j}\xi_{k,k,l}^{\top}Z_{l}^{\perp},$
while the third is a set of models for each mediator given covariates, the treatment, conditional on the treatment and all prior mediators, i.e., for all $k\in[K-1]$,
$\mathbb{E}[M_{k+1}\mid X,\overline{Z}_{k},\overline{1}_{k+1}]=$ $\theta_{k,0}+\delta_{k,1}^{T}X^{\perp}+\sum_{j=1}^{k}\delta_{k,j+1}^{T}Z_{j}^{\perp}.$ (12)
These models differ from conventional linear regression in that (i) pre-treatment variables are centered around their marginal means, and (ii) post-treatment confounders $Z_{k}\forall k\in\{1,\dots K\}$ are centered around their conditional means given all antecedent variables. Under Assumptions 3-5 in the main text, and assuming that the outcome and mediators are linear in pre- and post-treatment confounders, the treatment, and prior mediators, and that all necessary interaction terms have been accounted for, then the ATE $\tau_{0}$ can be obtained from the linear model $\mathbb{E}[Y\mid X,A]$ as $\lambda_{1}$, and coefficients from the models $\mathbb{E}[Y\mid X,A,\overline{Z}_{k},\overline{M}_{k}]$ and $\mathbb{E}[M_{k+1}\mid X,A,\overline{Z}_{k},\overline{M}_{k}]$ yield estimates of the components of the decomposition as follows:
$\tau_{k}$ $=\mathbb{E}[Y(\overline{1}_{k+1})-Y(\overline{1}_{k},0)]=\beta_{k,k},\forall k\in\{1,\dots,K\},$
$\Delta_{k}$ $=\mathbb{E}[Y(\overline{1}_{k+1},\underline{0}_{k+2})-Y(\overline{1}_{k},\underline{0}_{k+1})]=\beta_{k+1,k-1},\forall k\in\{0,\dots,K-1\},$
$\pi_{k+1}$ $=\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]=\theta_{k,0},\forall k\in\{0,\dots,K-1\}.$
I state the RWR estimation procedure formally in the following algorithm:
RWR.
  1. 1.
    For each of the baseline confounders, compute $\hat{X}^{\perp}=X-\mathbb{P}_{n}[X]$, where $\mathbb{P}_{n}[\cdot]$ denotes empirical average.
  2. 2.
    Fit $\mathbb{\hat{E}}[Y\mid X,A]$ using the linear specification shown above; an estimate of $\tau_{0}$ is given by $\hat{\lambda_{1}}$.
  3. 3.
    For each set of post-treatment confounders $Z_{k}$, $k\in\{1,\dots,K\}$, compute $Z_{k}^{\perp}=M_{k-1}\big{[}Z_{k}-\mathbb{E}[Z_{k}\mid X,\overline{Z}_{k-1},M_{k-1}=1]\big{]}$ where an overbar denotes a vector of variables such that $\overline{Z}_{k}=(Z_{1},\dots,Z_{k})$, by fitting a regression of $Z_{k}$ on $X$ and $\overline{Z}_{k-1}$ among units with $M_{k-1}=1$ and then calculating the residuals.
  4. 4.
    For each $k\in\{1,\dots K\}$:
    1. (a)
      compute least squares estimates of equations 11 and 12, using estimates of $X^{\perp}$ and $Z_{k}^{\perp}$.
    2. (b)
      compute $\text{$\hat{\tau}$}_{k}^{\text{RWR}}=\hat{\beta}_{k,k}$, $\hat{\Delta}_{k-1}^{\text{RWR}}=\hat{\beta}_{k,k-1}$, and $\hat{\pi}_{k}^{\text{RWR}}=\hat{\theta}_{k-1,0}$.
  5. 5.
    Compute the decomposition using $\hat{\tau}_{k}$, $\hat{\Delta}_{k}$ and $\hat{\pi}_{k+1}$, and estimating the covariance terms as $\hat{\eta}_{k}^{\text{RWR}}=\hat{\beta}_{k-1,k}-\hat{\beta}_{k,k-1}-\hat{\beta}_{k,1}\hat{\theta}_{k-1,k}$, and the continuation effects as $\hat{\theta}_{k}^{\text{RWR}}=(\Pi_{j=1}^{k}\hat{\pi}_{j}^{\text{RWR}})\hat{\Delta}_{k}^{\text{RWR}}+(\Pi_{j=1}^{k-1}\hat{\pi}_{j}^{\text{RWR}})\hat{\eta}_{k}^{\text{RWR}}$.
Standard errors and confidence intervals can then be obtained via the non-parametric bootstrap, or by using their asymptotic analytic variance. Specifically, let $\hat{\theta_{k}^{*}}\triangleq(\hat{\beta}_{k,0},\hat{\beta}_{k,1},\theta_{k,0})$ denote a set of parameters. Under the above models, we have that $\hat{\theta}^{*}=\{\hat{\lambda}_{1},\theta_{1}^{*},\dots\theta_{K}^{*}\}$ solves $\mathbb{P}_{n}[g(O;\hat{\theta}^{*})]=0$, where $g(O;\theta^{*})$ is the set of stacked moment conditions with solution $\hat{\theta}^{*}$. Under standard regularity conditions [Newey and McFadden, 1994], under correct specification of Models 9-11 wherein all residualized quantities are estimated via linear models, the set $\hat{\theta}^{*}$ is consistent and asymptotically normal, such that $\sqrt{n}\big{(}\hat{\theta}^{*}-\theta^{*})$ converges to a mean-zero normal distribution with finite variance $V=G^{-1}\Omega(G^{-1})^{\top}$, where $\Omega=\mathbb{E}[g(O;\theta^{*})g(O;\theta^{*})^{\top}]$, and where $G=\mathbb{E}[\frac{\partial g(O;\theta^{*})}{\partial\theta^{\top}}]$. It follows by a simple application of the Delta Method that the set $\hat{\gamma_{k}^{*}}\triangleq(\hat{\tau}_{k}^{\text{RWR}},\hat{\Delta}_{k-1}^{\text{RWR}},\hat{\pi}_{k}^{\text{RWR}},\hat{\eta}_{k}^{\text{RWR}},\hat{\theta}_{k}^{\text{RWR}})\forall k\in[K]$ is also consistent and asymptotically normal.
2 A simulation study
In this section, I evaluate the finite-sample performance of my two estimation procedures via a simulation experiment. Specifically, I compare how the DML estimator proposed in Section 3 (as well as the parametric, RWR estimator described in Appendix 1) perform under different degrees of misspecification. Without loss of generality, I focus on the single-mediator setting, and focus on the path-specific effect $A\to M\to Y$. I generate simulations of observed data $O=(X_{1},X_{2},A,Z,M,Y)$ as follows:
$U_{1},U_{2},U_{3},U_{4}$ $\sim\text{MVN}\big{(}0_{4},I_{4}\big{)}$
$X_{1}$ $\sim\text{N}\big{(}(U_{1},U_{2},U_{3},U_{4})\beta_{X_{1}},1\big{)}$
$X_{2}$ $\sim\text{N}\big{(}(U_{1},U_{2},U_{3},U_{4})\beta_{X_{2}},1\big{)}$
$A$ $\sim\text{Bern}\big{(}\text{\text{logit}}^{-1}\big{[}(1,X)\beta_{A}\big{]}\big{)}$
$Z|A=1$ $\sim\text{Bern}\big{(}\text{\text{logit}}^{-1}\big{[}(1,X)\beta_{Z}\big{]}\big{)}$
$M|A=1\sim\text{Bern}\big{(}\text{\text{logit}}^{-1}\big{[}(1,X,Z)\beta_{M}\big{]}\big{)}$ $\quad M|A=0=0$
$Y$ $\sim\text{N}\big{(}(1,A,X,AZ,AM)\beta_{Y},1\big{)}.$
The coefficients $(\beta_{X_{1}},\beta_{X_{2}},\beta_{Y})$ are drawn from a $\text{Unif}\big{(}-1,1\big{)}$ , while the coefficient $\beta_{A}$ is drawn from a $\text{Unif}\big{(}-0.5,0.5\big{)}$ distribution. In order to test how the DML and RWR methods perform when the relevant models are misspecified, I also construct transformations of the observed covariates ($X_{i}^{*},Z_{i}^{*}$) as follows, employing a similar setup to Kang and Schafer [2007]:
$X_{1}^{*}$ $=(\text{exp}(X_{1}/2)-1)^{2}$
[…]
For each simulated dataset, I construct two estimates of the path-specific effect $\theta_{1}$ ($A\to M\to Y$) by estimating the parameter set ($\tau_{0}$, $\Delta_{0},\Delta_{1},\pi_{1})$ via the RWR and DML procedures described in Section 3. Standard errors for the coverage rates are computed via the estimated variance of the estimated EIFs for the DML approach, and via the nonparametric bootstrap with 1000 replications for the RWR procedure. For the DML estimator, for each component involved in the decomposition, I construct a Neyman-orthogonal “signal” using its EIF. The recentered EIFs for each component are shown below:
$M^{*}(1)$ $=\gamma_{1}(X)+\frac{\mathbb{I}(A=1)}{\pi_{0}(X,1)}(M-\gamma_{1}(X)),$
$Y^{*}(a)$ $=\mu_{0}(X,a)+\frac{\mathbb{I}(A=a)}{\pi_{0}(X,a)}(Y-\mu_{0}(X,a)),\;\text{for }a\in\{0,1\}$
$Y^{*}(1,m_{1})$ $=\nu_{1}(X,m)+\frac{\mathbb{I}(A=1)\mathbb{I}(M=m)}{\pi_{0}(X,1)\pi_{1}(X,m_{1})}(Y-\mu_{1}(X,Z,m))$
$+\frac{\mathbb{I}(A=1)}{\pi_{0}(X,1)}(\mu_{1}(X,Z,m)-\nu_{1}(X,m)),\;\text{for }m\in\{0,1\}$
where
$\pi_{0}(X,a)$ $\triangleq\mathrm{Pr}[A=a\mid X]$
$\pi_{1}(X,m_{1})$ $\triangleq\mathrm{Pr}[M=m\mid X,A=1]$
$\gamma_{1}(X)$ $\triangleq\mathbb{E}[M\mid X,A=1]$
$\mu_{0}(X,a)$ $\triangleq\mathbb{E}[Y|X,A=a]$
$\mu_{1}(X,Z,m)$ $\triangleq\mathbb{E}[Y|X,A=1,Z,M=m]$
$\nu_{1}(X,m)$ $\triangleq\mathbb{E}[\mu_{1}(X,Z,m)|X,A=1].$
I run 1000 replications of this DGP and compute the average bias, the root mean square error (RMSE), and the coverage of nominal 95% confidence intervals for sample sizes of 250, 500, and 1000 and using either the ”correctly specified” covariates $\left(X_{1},X_{2},Z\right)$ and the ”incorrectly specified”, transformed versions $\left(X_{1}^{*},X_{2}^{*},Z^{*}\right)$. I calculate the true value of $\theta_{1}$ by recovering the true values of the parameter set ($\tau_{0}$, $\Delta_{0},\Delta_{1},\pi_{1})$ in each Monte Carlo simulation.
Figure 3 presents the results from this simulation exercise. Under correctly specified models, the DML and RWR estimators perform similarly, with each displaying low bias, low RMSE and close to nominal coverage at all sample sizes. Under incorrectly specified models, however, the DML and RWR approaches diverge in performance. Whereas the DML estimator under an incorrect feature space performs similarly to when it is used on the correct feature space, the RWR estimator performs much more poorly, displaying a large amount of bias that in fact grows with the sample size, a large RMSE, and coverage rates that are not close to nominal. In short, when the models are correctly specified, both the parametric and semiparametric approaches perform well; the strong performance of a semiparametric approach compared with a parametric estimation strategy becomes clearer under model misspecification.
Figure 3: Bias, RMSE, and coverage of DML and RWR estimators for $n=250,500,1000$. The left panel shows the performance of the DML and RWR estimators when the correct feature matrix is supplied to the estimators; the right panel shows the performance of the two estimators when an incorrect feature matrix is supplied to the estimators, as described in the main text.
3 Sensitivity analysis
How do my estimates of returns to different educational stages, as presented in the main text, tally with previous findings on the labor market returns to education? While previous work does not estimate quantities analogous to the direct and indirect effects of interest (i.e., the $\theta_{k}$ terms), some prior educational returns estimates are closely related to the net effect ($\tau_{k}$)) terms that inform the total, direct and indirect components of the decomposition. My estimate of the net effect of 4-year college enrollment ($\tau_{1}$) is large, at $54\%$, but not implausibly so. While Zimmerman [2014] and Smith et al. [2020] recover college earnings returns at around 2020by age 30 (exploiting admissions discontinuities in the Florida and Georgia state university systems), both of these studies estimate the earnings premium from attending a less selective 4-year college rather than a community college, for the marginally qualified university attendee. By contrast, $\tau_{1}$ captures the effect of 4-year college enrollment compared with community college and no college enrollment, pooling across the less selective colleges examined in Zimmerman [2014] and Smith et al. [2020], as well as over more selective colleges which could have greater earnings effects. Moreover, since $\tau_{1}$ represents an effect averaged over all individuals, it reflects a return among a broader population than the marginal college-goers examined in previous studies. [^4] My estimate of $\Delta_{1}$ (the direct effect of 4-year college attendance on earnings) further tallies with a similar quantity estimated by Scott-Clayton and Wen [2019]. On the intensive margin of employment (i.e. dropping respondents with zero observed earnings), the authors estimate a return to college attendance without degree completion of 0.21. While, theoretically, one might expect my estimate - which corresponds to the extensive employment margin (including respondents with zero observed earnings - to be larger, the fact that it is slightly smaller could reflect several factors, including the richer array of pre-college controls I use in my models, model mis-specification resulting from linearities imposed in prior work and, perhaps most importantly, collider-stratification biases induced by conditioning on BA completion in Scott-Clayton and Wen [2019]’s models (biases that are likely reduced by the inclusion of time-varying controls). As I discuss in the main text, an additional, especially point of comparison are instrumental variable (IV) estimates of returns to years of schooling. These results are in fact quite consistent with those I report in the main text.
Of course, an alternative explanation is that my estimates are upwardly biased by a large degree of unobserved confounding. While the sequential ignorability assumption facilitates identification of educational effect pathways under a weaker set of conditions than might be typically invoked in mediation settings, it is still strong and fundamentally unverifiable. To assess potential bias of the estimated MPSEs due to unobserved confounders not picked up in my covariate set $(X,\overline{Z}_{K})$, I propose a sensitivity analysis for each of the MPSEs.
Assume first that we have a binary unobserved confounder, $U$, for the treatment-outcome relationship. Assuming that $\alpha_{0}=\mathbb{E}[Y|x,a,U=1]-\mathbb{E}[Y|x,a,U=0]$ does not depend on $x$ or $a$, and further that $\beta_{0}=\text{Pr}[U=1|x,A=1]-\text{Pr}[U=1|x,A=0]$ does depend on $x$, for $\tau_{0}=\mathbb{E}[Y(1)-Y(0)]\triangleq\text{ATE}$, then $\text{bias}(\tau_{0})=\alpha\beta$ [VanderWeele and Arah, 2011],
Next, consider an unobserved binary confounder, $U_{k}$ that affects both $M_{k}$ and $Y$ for any $k\in\{1,\dots,K\}$. Then, under a weaker instantiation of Assumption 4 (Sequential Ignorability), i.e.,
$Y(\overline{1}_{k},m_{k})\perp\!\!\!\perp(A,\overline{M}_{k})|X,A,U_{k},\overline{Z}_{k},\overline{M}_{k-1}\forall k\in[K],$ (13)
which states that potential outcomes under an arbitrary transition sequence are independent of observed treatment and mediator values conditional on observed confounders $(X,\overline{Z}_{k})$ and unobserved confounders $U_{k}$. Under the following set of assumptions: (Assumption $A_{k}$) $\alpha_{k}=\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=1]-\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=0]$ does not depend on ($x,\overline{z}_{k},\overline{1}_{k},m_{k})$, and (Assumption $B_{k}$), $\beta_{k}=\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},m_{k}]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}$ does not depend on ($x,\overline{z}_{k}$), we can show that, for any $k\in\{1,\dots K\}$,
$\text{bias}(\tau_{k})$ $=\alpha_{k}\beta_{k},$
and, further, that
$\text{bias}(\Delta_{k-1})=-\alpha_{k}\beta_{k}\pi_{k},$
where $\pi_{k}=\int_{x}\int_{\overline{z}_{k}}\text{Pr}[M_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x)$, and is estimable from observed using the estimation strategies described previously. A contour plot showing bias-adjusted estimates of $\Delta_{k^{*}}$ and $\tau_{k}$ then enables assessment of how strong the unobserved confounder would need to be to reduce estimates of the direct and gross effects to zero. I illustrate these techniques in my empirical illustration below.
In order to assess the robustness of my empirical findings in the main text to potential violations of Assumption 4 (Sequential Ignorability), I implement this sensitivity analysis discussed above.
Figure 4 below displays a set of contour plots, which capture the bias-corrected estimates of the $(\Delta_{k},\tau_{k})$ terms under varying degrees of confounding (that is, under different values of $\alpha_{k}$ and $\beta_{k}$). For example, the level set marked “0” corresponds to values of $(\alpha_{k},\beta_{k})$ required in order for the unobserved confounder to fully “explain away” estimates $(\Delta_{k},\tau_{k})$ (i.e., to reduce their true values to zero). Importantly, each row corresponds to a different set of $(\alpha_{k},\beta_{k})$ terms for a given $U$, such that the top row corresponds to $(\alpha_{0},\beta_{0})$, while the second row corresponds to $(\alpha_{1},\beta_{1})$, and so on.
For simplicity, I consider $U$ to be an unmeasured binary confounder that is (marginally) positively associated with each transition $A,M_{1},\dots M_{3}$ as well as with adult earnings $Y$. To benchmark the hypothetical behavior of $U$, for each plot, I also display the values of $(\alpha_{k},\beta_{k})$ that would correspond to a $U$ that behaved similarly to a given confounder that I do observe in the data: an indicator for whether an individual’s test score on the ASVAB is above the median. In each plot, I mark this point and label it “Ability”. For each plot, I also mark the point on the zero contour that corresponds to $\alpha_{k}=\beta_{k}$ (i.e, the point at which the unobserved confounder’s associations with the treatment and with the outcome are equal, and reduce the true value of the parameter to zero).
I focus on estimates of $\tau_{0}$ and $\Delta_{0}$ to assess how robust my primary conclusion - that the ATE of high school completion is overwhelmingly mediated by high school’s direct effect on earnings - is to unobserved confounding. Bias-adjusted estimates of the ATE $\tau_{0}$ are presented in the top row of Figure 4. Since $U$ is assumed to be positively associated with both $A$ and $Y$, $\tau_{0}$ is overestimated and suffers from a bias of $\alpha_{0}\beta_{0}$. My estimate of $\tau_{0}$ at $0.67$ is nevertheless quite robust: if $U$ had similar effects to ability, the effect would be reduced by $0.06$ log points, to $0.61$, still implying a high earnings premium to high school completion overall in excess of $84\%$.
How do my estimates of the direct effect of high school completion $\Delta_{0}$ (and, in particular, about the proportion of the total effect that is direct) fare under unobserved confounding? The second row of Figure 4 considers bias-adjusted estimates of $\Delta_{0}=\theta_{0}$ under different values of $(\alpha_{1},\beta_{1})$, which correspond to the effects of an unobserved confounder $U$ (marginally) positively associated with both $M_{1}$ and with $Y$. As described above, in this scenario, $\Delta_{0}$ is affected by a bias of $\alpha_{1}\beta_{1}\pi_{1}$. Importantly, even if the unobserved confounder $U$ is marginally positively associated with high school graduation ($A$), the conditional association between $U$ and $A$ may be zero or even negative since $M_{1}$ is a collider of $A$ and $U$. In the case that the conditional association between $U$ and $A$ is negative, $\alpha_{1}\beta_{1}\pi_{1}$ would be positive, implying an overestimation of the direct effect $\theta_{0}$. On the plot, I show estimates of $U$ if it behaved similarly to the ability variable. Indeed, despite the fact that ability is marginally positively associated with high school completion (top row of Figure 4), its conditional association - conditional on college attendance - is depressed to zero. Thus, it would take an extreme form of confounding for $\theta_{0}$ to be largely different from its estimated value of .46.46. In this way, my primary finding that the ATE of high school graduation is overwhelmingly mediated via its direct effect remains highly robust to patterns of unobserved confounding, under my set of simplifying assumptions.
Figure 4: Sensitivity Analysis for the “gross effect” $(\tau_{k})$ and “direct effect” $(\Delta_{k})$ terms in decomposition. Each row corresponds to a different set of ($\alpha_{k},\beta_{k})$ terms, where $\alpha_{k}=\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=1]-\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=0]$ parameterizes the effect of $U_{k}$ on Y, and $\beta_{k}=\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},m_{k}]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}$ parameterizes the effect of $M_{k}$ on $U_{k}$. Each row corresponds to a different set of $(\alpha_{k},\beta_{k})$ terms. For example, the top row corresponds to $(\alpha_{0},\beta_{0})$, while the second row corresponds to $(\alpha_{1},\beta_{1})$, and so on.
4 Further details on variable construction and education groups
Variable construction
In an effort to satisfy the sequential ignorability assumption (Assumption 4), I include a large array of covariates in my models for the effects of completing educational transitions on labor market outcomes. Figure 5 summarizes my assumed data-generating process for the empirical example. In addition to including information on respondent demographics (gender, race, ethnicity, age at 1997), and observed pre-college performance such as overall high school GPA and test score on the Armed Services Vocational Aptitude Battery (ASVAB), I include detailed information on socioeconomic background (parental education, parental income, parental asset, co-residence with both biological parents, presence of a paternal figure, rural residence, southern residence), an index of substance use, an index of delinquency, whether the respondent had any children by age 18), and peer and school-level characteristics (measures of peers’ college expectations and behaviors). Both parental income and parental asset variables are transformed to 2023 dollars.
[…]
How convincingly do I satisfy the sequential ignorability assumption? Despite the inclusion of a comprehensive set of background covariates in my models, it is possible that observed variables do not perfectly proxy for all important confounders jointly affecting education and earnings. In particular, researchers often argue that important variables, such as students’ innate ability, ambition, and detailed forms of socioeconomic advantage, confound observational estimates of educational returns [e.g. Carneiro et al., 2011]. While some research suggests that observational estimates of earnings returns may well capture actual returns to education and that the degree of observational bias may be rather small [Card, 1999], it is of course impossible to quantify the true extent of the bias in the estimates I produce. The sensitivity analysis described above provides a step towards this goal.
I note that my assumption of ignorability of $M_{3}$ without conditioning on intermediate variables $Z_{3}$ is perhaps the strongest assumption I make. For example, many individuals take time off to work before enrolling in graduate school, and labor market experience and earnings gained in the interim period between college completion and graduate school enrollment may confound the latter variable’s effects on earnings. Nevertheless, including a measure of labor market characteristics for this period is difficult because some respondents enroll directly in graduate school after BA completion, such that pre-graduate school earnings variables would be undefined for these individuals.
Table 2 shows conditional means of respondent attributes $X$ and $Z$ for the full (imputed) and restricted (non-imputed) samples, showing first the mean among the full population of high school goers, and progressively restricting the sample from (i) high school (HS) non-completers, to (ii) HS graduates, to (iii) college attendees and, finally, to (iv) BA completers. Imputed and non-imputed means - shown without and with brackets, respectively - are highly similar across variables. As I progressively restrict the sample to those who attained higher educational levels, variables capturing components of socioeconomic advantage (such as parental income, parental education and household net worth) increase monotonically in value. Background covariates measuring aspects of the school environment (such as peers’ college expectations - which is an indicator for whether over $90$ of a respondent’s peers expected to go to college) behave similarly. I also see that students who progress to higher educational levels have higher levels of pre-college ability: HS non-completers have on average an ASVAB Percentile score of 22.3, compared with only in excess of 70 among BA completers. Similarly, college-goers average high school GPA is approximately .5.5 higher than high school graduates overall (regardless of whether or not they proceed to college). Nevertheless, the association between high school GPA and attainment declines at higher educational levels: BA completers have only on average a .11.11 higher a high school GPA than the pooled group of college goers, irrespective of their BA completion status. At this stage, college GPA appears to matter more: college goers overall have on average a college GPA of $2.77$, while BA completers’ average college GPA is $3.07$.
Educational groups: raw mean earnings
Table 3 (column 2) presents the proportion of individuals who have attained each level of education constructed above. By age 22, a small, but not insignificant, proportion of individuals who enroll in high school do not complete their studies ($13$), and by this same age, just over $40\%$ of individuals have attended a 4-year college. By age 29, $29$ of individuals have attained a Bachelor’s degree or higher. These estimates of high school completion and BA completion align closely both with those reported in previous studies that employ the NLSY97 (e.g. Scott-Clayton and Wen [2019], as well as with those reported in the Current Population Survey (CPS). Table 3 (columns 3-4) also shows mean log earnings by educational group (column 3), alongside the estimated gap between these means and mean log earnings among high school non-completers (column 4). High school dropouts earn an average of 9.07 log earnings, while groups with higher levels of attainment earn successively more than high school dropouts, though at a decreasing rate. High school graduates earn on average $1.11$ log earnings more than high school non-completers, implying an earnings premium in excess of $200\%$ ($\exp(1.11)-1)$, while college goers earn on average $1.5$ log earnings more than high school non-completers (or 0.6 log earnings more than high school graduates). At the highest end, graduate school goers earn on average $10.88$ log earnings. These educational premia are extremely high, since they reflect both the causal effect of a given educational level as well as the effects of individual, geographic and family factors correlated both with attainment and with adult earnings. To net out these patterns of selection, we need to turn to estimates of the MPSE decomposition, as well as its constituent components.
Figure 5: DAG showing the hypothesized causal relationships between high school completion $A$ and adult earnings $Y$ via mediators $M_{1}$, $M_{2}$ and $M_{3}$.
Table 2: Conditional means of background and college-level attributes by sample type (imputed dataset and non-imputed dataset).
Full Population HS Non-Completers HS Graduates College Goers BA Completers
Female 0.50 (0.50) 0.44 (0.47) 0.51 (0.51) 0.55 (0.56) 0.57 (0.58)
Black 0.17 (0.16) 0.25 (0.24) 0.16 (0.15) 0.13 (0.10) 0.10 (0.09)
Hispanic 0.14 (0.12) 0.20 (0.18) 0.13 (0.12) 0.09 (0.08) 0.08 (0.07)
Parental Income 80,235 (79,243) 43,308 (43,863) 85,898 (83,865) 110,240 (109,636) 118,026 (114,329)
Parental Education 12.85 (12.84) 11.18 (11.24) 13.11 (13.04) 14.06 (14.13) 14.38 (14.34)
Household Net Worth 175,513 (176,097) 61,997 (62,318) 192,920 (190,959) 272,606 (284,029) 306,177 (307,101)
Lived w Biological Parents 0.52 (0.51) 0.30 (0.28) 0.55 (0.55) 0.67 (0.69) 0.72 (0.73)
Father Figure Present 0.75 (0.75) 0.61 (0.61) 0.77 (0.76) 0.84 (0.83) 0.86 (0.86)
Lived in Rural Area 0.28 (0.30) 0.25 (0.27) 0.28 (0.31) 0.28 (0.32) 0.28 (0.31)
Lived in South 0.36 (0.35) 0.44 (0.42) 0.35 (0.34) 0.33 (0.30) 0.31 (0.29)
Children by 18 0.07 (0.07) 0.21 (0.24) 0.05 (0.05) 0.01 (0.01) 0.01 (0.01)
Substance Abuse Score 1.09 (1.10) 1.32 (1.37) 1.06 (1.07) 0.86 (0.82) 0.82 (0.77)
Delinquency Score 1.38 (1.39) 2.07 (2.13) 1.27 (1.29) 0.91 (0.88) 0.81 (0.79)
Peers’ College Expectations 0.56 (0.56) 0.40 (0.38) 0.59 (0.58) 0.69 (0.69) 0.72 (0.71)
Property Stolen at School 0.24 (0.23) 0.28 (0.29) 0.23 (0.22) 0.21 (0.20) 0.20 (0.19)
Threatened at School 0.22 (0.23) 0.30 (0.33) 0.20 (0.21) 0.14 (0.13) 0.13 (0.11)
In a Fight at School 0.16 (0.16) 0.32 (0.34) 0.14 (0.13) 0.07 (0.07) 0.06 (0.06)
ASVAB Percentile 48.36 (48.52) 22.27 (21.77) 52.37 (52.02) 67.22 (69.66) 70.66 (71.82)
High School GPA 2.84 (2.81) 2.12 (2.14) 2.95 (2.90) 3.30 (3.33) 3.41 (3.43)
Stem Major 0.17 (0.19) 0.18 (0.19)
College GPA 2.78 (2.86) 3.07 (3.12)
Earnings (\$) 46,505 (45,151) 21,597 (21,566) 46,505 (45,151) 64,984 (64,595) 72,374 (70,212)
Log ( Earnings + c ) 10.03 (10.06) 9.07 (9.14) 10.03 (10.06) 10.57 (10.65) 10.73 (10.78)
Note: Numbers denote means for the imputed sample (non-imputed sample). Means are adjusted for multiple imputation via Rubin’s (1987) method, and all statistics are calculated using NLSY97 sampling weights.
Table 3: Means of observed log earnings by educational participation, and earnings gaps (versus high school non-completers).
Group Population Proportion Log Earnings Gap (vs HS Non-Completers)
HS Non-Completers 0.13 9.07 (0.05)
HS Graduates 0.87 10.18 (0.02) 1.11 (0.05)
College Goers 0.41 10.57 (0.03) 1.5 (0.05)
BA Completers 0.29 10.73 (0.03) 1.66 (0.06)
Grad. School Goers 0.09 10.88 (0.05) 1.81 (0.07)
Note: The category “High School Non-Completers” captures all individuals who attended high school but did not obtain a high school diploma; “High School Graduates” refers to those individuals who graduated high school, regardless of their subsequent educational experiences (i.e., whether or not they proceeded to college); “College Goers” refers to individuals who attended a 4-year college, irrespective of whether they completed their degree; “BA Completers” denotes individuals who completed a Bachelor’s degree, while “Grad. School Goers” captures individuals who participated in a graduate-level degree program. A small constant of \$1,000\\$1,000 is added to observed earnings before taking the log. All statistics are computed with a monotonicity assumption imposed on the observed data (i.e. such that all individuals who complete a given educational level are coded as having completed all prior levels). All statistics are calculated using NLSY97 sampling weights, and standard errors are in parentheses.
5 Results without imputation of missing covariates
In the main text, I report estimates of the MPSE decomposition for the ATE of high school attendance on logged annual earnings for the full sample of NLSY97 respondents with non-missing educational information and non-missing earnings ($N=7,305$). A very large number (approximately $50\%$) of these respondents are missing information on one or more of the covariates $(X,Z)$ used in the models in order to identify the decomposition components, due especially to the large amount of missingness on variables such as parental assets ($25\%$ of the sample), and the Armed Services Vocational Aptitude Battery (ASVAB) test ($20\%$ of the sample). To assess the sensitivity of my primary conclusions to the use of multiple imputation as opposed to dropping observations with missing values, I replicate my DML and RWR estimates on a non-imputed analytic sample ($N=3,735)$. Figure 6 below shows the results of this exercise. For both estimation procedures, results under multiple imputation and non-imputation are highly similar. As is to be expected, imputation reduces standard errors significantly, especially for the parametric RWR procedure. Further, the greatest variability between imputed and non-imputed results come from effects pertaining to high school completion, perhaps because patterns of missingness are correlated with educational attainment. Despite this, because the total effect $\tau_{0}$ and direct effect $\theta_{0}$ are similarly attenuated in the imputed sample, the overall conclusion about the importance of the direct effect in explaining the ATE remains unaffected.
Figure 6: Decomposition of the Average Total Effect (ATE) of High School Graduation on Logged Earnings Under Multiple Imputation (MI) and Under Dropping Observations with Missing ($X,Z$) Values. Results with multiple imputation (purple lines) are reproduced from the main text ($N=7,305$); results without multiple imputation (red lines) employ a sample restricted to respondents with observed values for all covariates used ($N=3,735$).
6 Results under alternative definitions of earnings
In the main text, I report estimates of the MPSE decomposition components for the ATE of high school attendance on logged annual earnings. Logged annual earnings in the main text are defined as the log of observed annual earnings plus a small constant of \$1,000\\$1,000, in order to accommodate respondents with zero observed annual earnings. In order to assess the sensitivity of the reported results to the choice of this constant, I replicate the main analyses under alternative definitions of earnings. Figure 7 reports estimates of the direct and indirect effects under a series of different constants $c$ added to pre-logged annual earnings, for $c\in\{10,100,1000\}$, while Figure 8 shows estimates of these direct and indirect effects for observed annual earnings in dollar values. Beginning with Figure 7, I see that, while for the indirect effects $\theta_{1}$, $\theta_{2},$ and $\theta_{3}$, both DML and RWR estimates are quite consistent under these different constants, estimates of the total effect $\tau_{0}$ as well as the direct effect $\Delta_{0}$ are quite sensitive to the choice of constant. Specifically, lower constant values correspond with large increases in the DML estimate of $\tau_{0}$ from $0.67$ ($c=1000$) to $1.20$ ($c=10$), and of $\Delta_{0}$ from $0.47$ ($c=1000$) to $0.89$ ($c=10$). This is because individuals with less than a high school degree are more likely than their higher-educated counterparts to have zero or low earnings, making their logged earnings rather sensitive to the choice of constant. Nevertheless, because the total effect $\tau_{0}$ and direct effect $\theta_{0}$ are similarly affected by the change in constant value, the importance of the direct effect in explaining the ATE of high school completion is reinforced. In particular, the proportion of the total effect that is direct is estimated to be $70\%$, $74\%$ and $72\%$ under each of $c=10,100,1000$. Turning next to Figure 7, under DML, estimates of high school completion increases earnings in expectation by roughly \$16,500\\$16,500, corresponding to an earnings return of approximately $53$ relative to a baseline of \$31,300\\$31,300 without high school graduation ($\mathbb{E}[Y(0)]$) . Almost half ($\frac{7824}{16454.}\cdot 100=47.5\%$) of the total effect is estimated to operate directly, with $29\%$ and $22\%$ mediated via college attendance and college completion, respectively.
Figure 7: Decomposition of the Average Total Effect (ATE) of High School Graduation on Logged Earnings Under Alternative Definitions of Earnings. The figure shows estimates of the total effect ($\tau_{0}$) as well as the indirect effects $\Delta_{0},\Delta_{1},\dots\dots,\Delta_{K}$ when constants of 10, 100 and 1000, respectively, are added to raw annual earnings (in dollar amounts) before taking the log.
Figure 8: Decomposition of the Average Total Effect (ATE) of High School Graduation on Logged Earnings with Different Definitions of Earnings. The figure shows estimates of the total effect ($\tau_{0}$) as well as the indirect effects $\Delta_{0},\Delta_{1},\dots\dots,\Delta_{K}$ when constants of 10, 100 and 1000 are added to raw annual earnings (in dollar amounts) before taking the log.
7 Extension to multivalued, discrete mediators
The main text considers a decomposition of the ATE in the case of binary monotonic mediators (i.e., educational transitions), but the framework can straightforwardly be extended to accommodate categorical transitions. Such a decomposition is especially appealing when we consider the variegated and complex trajectories that individuals take through the US postsecondary system, that dichotomizing transitions invariably misses. In particular, in the early 2010s, under $40\%$ of high school graduates immediately enrolled in a four-year college, with around $30\%$ immediately attending a two-year college. Approximately one third of immediate two-year college attendees will then progress to a four-year college; indeed, nearly $50\%$ of BA recipients previously attended a public two-year college in their educational careers. To illustrate the MPSE decomposition for multivalued, discrete mediators in the case of a single mediator, consider high school graduation $A$, and immediate college enrollment $M\in\{\text{none},\text{2yr},\text{4yr}\}$ denotes whether an individual did not pursue postsecondary education, or instead attended a 2-year or 4-year college. Assumption 2 is similarly assumed to hold in this context, such that $M(0)=0$ (high school non-completers cannot pursue any form of postsecondary education). Under this assumption, I obtain the following multivariate decomposition:
ATE $=\mathbb{E}[Y(1)-Y(0)]$
$=\mathbb{E}[Y(1,0)-Y(0)]$
$+\mathbb{P}[M(1)=\text{2yr}]\mathbb{E}[Y(1,\text{2yr})-Y(1,0)]+\text{cov}\big{[}\mathbb{I}[M_{1}(1)=\text{2yr}],Y(1,\text{2yr})-Y(1,0)\big{]}$ (14)
$+\mathbb{P}[M(1)=\text{4yr}]\mathbb{E}[Y(1,\text{4yr})-Y(1,0)]+\text{cov}\big{[}\mathbb{I}[M_{1}(1)=\text{4yr}],Y(1,\text{4yr})-Y(1,0)\big{]}.$ (15)
Each of these quantities can be identified under Assumptions 3-5 above. Further, Equations 14 and 15 can be further decomposed into the direct and gross effects of each of these transitions in a way analogous to Equation 4 (for example, the gross effect of 2-year attendance can be decomposed into its direct effects, net of subsequent transfer to a 4-year college, and its indirect effect via 4-year transfer, which itself can be further decomposed into direct and gross effects).
8 Description of EIFs used in empirical illustration
For each component involved in the MPSE, I construct a Neyman-orthogonal “signal” using its EIF, whose exact form depends on whether each set of intermediate confounders is empty or not. Figure 5 in the main text shows a potential data-generating process for the direct and indirect (continuation) effects of high school graduation on adult earnings, via three transitions: college attendance ($M_{1}$), BA completion ($M_{2}$), and graduate school attendance ($M_{3}$). I assume that a set of pre-college characteristics serve as confounders for the $A-(M_{1},M_{2},M_{3},Y)$ relationships, and that a set of post-secondary confounders $Z$ confound the $M_{2}-(M_{3},Y)$ relationships.
Under these assumptions for the various sets of confounders, my MPSE decomposition implies that, in the case of the four transitions (one treatment and three mediators), it suffices to estimate the following three sets of parameters: (i) four direct effects $\Delta_{k},k\in[0\dots,3]$, where $\Delta_{3}=\tau_{3}$, (ii) four gross effects $\tau_{k},k\in[0\dots,3]$, where $\tau_{0}=\mathrm{ATE}$, (iii) three mediator terms, $\pi_{k},k\in[1\dots,3]$. All components in the three-mediator decomposition can then be estimated as functions of these parameters. For each of these target parameters, I construct a Neyman-orthogonal signal using its efficient influence function. Because of my assumed data-generating process, which maintains that there is only a single set of intermediate confounders (as opposed to a separate set of confounders for each mediator), the EIF for each estimand involved in the decomposition simplifies somewhat. Specifically, the recentered EIFs for each component in the decomposition are shown below:
$M_{1}^{*}(1)$ $=\gamma_{1}(X)+\frac{\mathbb{I}(A=1)}{\pi_{0}(X,1)}(M_{1}-\gamma_{1}(X)),$
$M_{2}^{*}(1,1)$ $=\gamma_{2}(X)+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)}{\pi_{0}(X,1)\pi_{1}(X,1)}(M_{2}-\gamma_{2}(X)),$
$M_{3}^{*}(1,1,1)$ $=\mathbb{E}[\gamma_{3}(X,Z)|X,A=1,M_{1}=1]$
$+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)}{\pi_{0}(X,1)\pi_{1}(X,1)}(\gamma_{3}(X,Z)-\mathbb{E}[\gamma_{3}(X,Z)|X,A=1,M_{1}=1])$
$+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)\mathbb{I}(M_{2}=1)}{\pi_{0}(X,1)\pi_{1}(X,1)\pi_{2}(X,Z,1)}(M_{3}-\gamma_{3}(X,Z)),$
$Y^{*}(a)$ $=\mu_{0}(X,a)+\frac{\mathbb{I}(A=a)}{\pi_{0}(X,a)}(Y-\mu_{0}(X,a)),\;\text{for }a\in\{0,1\}$
$Y^{*}(1,m_{1})$ $=\mu_{1}(X,m_{1})+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=m_{1})}{\pi_{0}(X,1)\pi_{1}(X,m_{1})}(Y-\mu_{1}(X,m_{1})),\;\text{for }m_{1}\in\{0,1\}$
$Y^{*}(1,1,m_{2})$ $=\mathbb{E}[\mu_{2}(X,Z,m_{2})|X,A=1,M_{1}=1]$
$+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)}{\pi_{0}(X,1)\pi_{1}(X,1)}(\mu_{2}(X,Z,m_{2})-\mathbb{E}[\mu_{2}(X,Z,m_{2})|X,A=1,M_{1}=1])$
$+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)\mathbb{I}(M_{2}=m_{2})}{\pi_{0}(X,1)\pi_{1}(X,1)\pi_{2}(X,Z,m_{2})}(Y-\mu_{2}(X,Z,m_{2})),\;\text{for }m_{2}\in\{0,1\}$
$Y^{*}(1,1,1,m_{3})=$ $\mathbb{E}[\mu_{3}(X,Z,m_{3})|X,A=1,M_{1}=1]$
$+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)}{\pi_{0}(X,1)\pi_{1}(X,1)}\left(\mu_{3}(X,Z,m_{3})-\mathbb{E}[\mu_{3}(X,Z,m_{3})|X,A=1,M_{1}=1]\right)$
$+\frac{\mathbb{I}(A=1)\mathbb{I}(M_{1}=1)\mathbb{I}(M_{2}=1)\mathbb{I}(M_{3}=m_{3})}{\pi_{0}(X,1)\pi_{1}(X,1)\pi_{2}(X,Z,1)\pi_{3}(X,Z,m_{3})}\left(Y-\mu_{3}(X,Z,m_{3})\right)\;\text{for }m_{3}\in\{0,1\},$
where
$\pi_{0}(X,a)$ $\triangleq\mathrm{Pr}[A=a\mid X]$
$\pi_{1}(X,m_{1})$ $\triangleq\mathrm{Pr}[M_{1}=m_{1}\mid X,A=1]$
$\pi_{2}(X,Z,m_{2})$ $\triangleq\mathrm{Pr}[M_{2}=m_{2}\mid X,A=1,M_{1}=1,Z]$
$\pi_{3}(X,Z,m_{3})$ $\triangleq\mathrm{Pr}[M_{3}=m_{3}\mid X,A=1,M_{1}=1,Z,M_{2}=1]$
$\gamma_{1}(X)$ $\triangleq\mathbb{E}[M_{1}\mid X,A=1]$
$\gamma_{2}(X)$ $\triangleq\mathbb{E}[M_{2}\mid X,A=1,M_{1}=1]$
$\gamma_{3}(X,Z)$ $\triangleq\mathbb{E}[M_{3}\mid X,A=1,M_{1}=1,Z,M_{2}=1]$
$\mu_{0}(X,a)$ $\triangleq\mathbb{E}[Y|X,A=a]$
$\mu_{1}(X,m_{1})$ $\triangleq\mathbb{E}[Y|X,A=1,M_{1}=m_{1}]$
$\mu_{2}(X,Z,m_{2})$ $\triangleq\mathbb{E}[Y|X,A=1,M_{1}=1,Z,M_{2}=m_{2}]$
$\mu_{3}(X,Z,m_{3})$ $\triangleq\mathbb{E}[Y|X,A=1,M_{1}=1,Z,M_{2}=1,M_{3}=m_{3}].$
9 Proofs and technical details
9.1 EIFs for $\eta_{k}$ and $\theta_{k}$ terms (Proposition 1)
Under Assumptions 3-5, the covariance component $\eta_{k}$ is identified as $\eta_{k}=\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k}$. Following [Kennedy, 2022, , p. 15], I let $\mathbb{IF}:\Psi\rightarrow L_{2}(\mathbb{P})$ denote the operator mapping the functionals $\{\Delta_{k},\pi_{k},\eta_{k}\}:\mathcal{P}\rightarrow\mathbb{R},$ $\forall\in[K]$ to their respective influence functions under the nonparametric model $\mathcal{P}$. First, by linearity of the EIF, $\mathbb{IF}(\eta_{k})$ is given by
$\mathbb{IF}(\eta_{k})=\mathbb{IF}(\tau_{k-1})-\mathbb{IF}(\Delta_{k-1})-\mathbb{IF}(\pi_{k}\tau_{k}).$
Since $\mathbb{IF}(\pi_{k}\tau_{k})$ can be written as follows $\mathbb{IF}(\pi_{k}\tau_{k})=\tau_{k}\mathbb{IF}(\pi_{k})+\pi_{k}\mathbb{IF}(\tau_{k})$, $\mathbb{IF}(\eta_{k})$ can be written as
$\mathbb{IF}(\eta_{k})$ $=\mathbb{IF}(\tau_{k-1})-\mathbb{IF}(\Delta_{k-1})-\big{(}\tau_{k}\mathbb{IF}(\pi_{k})+\pi_{k}\mathbb{IF}(\tau_{k})\big{)}$
$=\mathbb{IF}(\tau_{k-1})-\mathbb{IF}(\Delta_{k-1})-\tau_{k}\mathbb{IF}(\pi_{k})-\pi_{k}\mathbb{IF}(\tau_{k}).$
Noticing that we can rewrite this expression as
$\mathbb{IF}(\eta_{k})$ $=\mathbb{RIF}(\tau_{k-1})-\tau_{k-1}-\mathbb{RIF}(\Delta_{k-1})+\Delta_{k-1}-\tau_{k}\mathbb{RIF}(\pi_{k})+\tau_{k}\pi_{k}-\pi_{k}\mathbb{RIF}(\tau_{k})+\pi_{k}\tau_{k}$
$=\mathbb{RIF}(\tau_{k-1})-\mathbb{RIF}(\Delta_{k-1})-\tau_{k}\mathbb{RIF}(\pi_{k})-\pi_{k}\mathbb{RIF}(\tau_{k})+\pi_{k}\tau_{k}-\eta_{k},$
where $\mathbb{RIF}(\phi)=\mathbb{IF}(\phi)+\phi$, we can obtain the corresponding EIF-based estimator for $\eta_{k}$ by solving the empirical moment condition implied by setting the average of the above equation equal to 0, and plugging in the set of estimated nuisance functions:
[…]
Turning next to the influence functions for the continuation effects $\theta_{k}$, $k\in\{1,\dots K\}$, following the same logic as the above, we can write the EIF of $\theta_{k}$, $\mathbb{IF}(\theta_{k})$, as
$\mathbb{IF}(\theta_{k})=\mathbb{IF}(\Delta_{k})\prod_{j=1}^{k}\pi_{j}+\Delta_{k}\sum_{j=1}^{k}\mathbb{IF}(\pi_{j})\prod_{l:l\neq j}^{k}\pi_{l}+\mathbb{IF}(\eta_{k})\prod_{j=1}^{k-1}\pi_{j}+\eta_{k}\sum_{j=1}^{k-1}\mathbb{IF}(\pi_{j})\prod_{l:l\neq j}^{k-1}\pi_{l}.$
Rewriting this expression as
$\mathbb{IF}(\theta_{k})$ $=\mathbb{RIF}(\Delta_{k})\prod_{j=1}^{k}\pi_{j}+\Delta_{k}\sum_{j=1}^{k}\mathbb{RIF}(\pi_{j})\prod_{l:l\neq j}^{k}\pi_{l}+\mathbb{RIF}(\eta_{k})\prod_{j=1}^{k-1}\pi_{j}+\eta_{k}\sum_{j=1}^{k-1}\mathbb{RIF}(\pi_{j})\prod_{l:l\neq j}^{k-1}\pi_{l}$
$\;-k\Delta_{k}\prod_{j=1}^{k}\pi_{j}-(k-1)\eta_{k}\prod_{j=1}^{k-1}\pi_{j}-\theta_{k},$
we obtain the corresponding EIF-based estimator for $\theta_{k}$ as:
$\hat{\theta}_{k}^{\text{eif}}$ $=\widehat{\mathbb{RIF}}(\Delta_{k})\prod_{j=1}^{k}\hat{\pi}_{j}+\hat{\Delta}_{k}\sum_{j=1}^{k}\hat{\mathbb{RIF}}(\pi_{j})\prod_{l:l\neq j}^{k}\hat{\pi}_{l}+\widehat{\mathbb{RIF}}(\eta_{k})\prod_{j=1}^{k-1}\hat{\pi}_{j}+\hat{\eta}_{k}\widehat{\mathbb{RIF}}(\pi_{j})\prod_{l:l\neq j}^{k-1}\hat{\pi}_{l}$
$\;-k\hat{\Delta}_{k}\prod_{j=1}^{k}\hat{\pi}_{j}-(k-1)\hat{\eta}_{k}\prod_{j=1}^{k-1}\hat{\pi}_{j}.$
9.2 Semiparametric efficiency (Proposition 2)
[…]
$\hat{\varphi}_{km_{k}}^{\text{EIF}}-\varphi_{km_{k}}$ $=\mathbb{P}_{n}[m(O;\hat{\eta})]-P[m(O;\eta)]$
$=\mathbb{P}_{n}[m(O;\eta)]+\underbrace{P[m(O;\hat{\eta})-m(O;\eta)]}_{\triangleq R_{2}(\hat{\eta})}+\left(\mathbb{P}_{n}-P\right)[m(O;\hat{\eta})-m(O;\eta)],$ (16)
[…]
$P[m(O;\eta)]$ $=P\bigg{[}\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\frac{\mathbb{I}(M_{k}=m_{k})}{\pi_{km_{k}}\big{(}X,\overline{Z}_{k}\big{)}}\prod_{j=1}^{k-1}\frac{M_{j}}{\pi_{j1}\big{(}X,\overline{Z}_{j}\big{)}}\right)\left(Y-\mu_{km_{k}}^{k}\left(X,\bar{Z}_{k}\right)\right)$
$\quad+\sum_{j=1}^{k}\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\prod_{l=1}^{j-1}\frac{M_{l}}{\pi_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\left(\mu_{jm_{k}}^{k}\left(X,\overline{Z}_{k}\right)-\mu_{j-1m_{k}}^{k}\left(X,\bar{Z}_{j-1}\right)\right)+\mu_{0}(X)\bigg{]}$
$=P\bigg{[}\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\frac{\mathbb{I}(M_{k}=m_{k})}{\pi_{km_{k}}\big{(}X,\overline{Z}_{k}\big{)}}\prod_{j=1}^{k-1}\frac{M_{j}}{\pi_{j1}\big{(}X,\overline{Z}_{j}\big{)}}\right)\cdot$
$\quad\left(\underbrace{\mathbb{E}[Y-\mu_{km_{k}}^{k}\left(X,\bar{Z}_{k}\right)|X,\bar{Z}_{k},\overline{1}_{k},m_{k}]}_{=0}\right)$
$\quad+\sum_{j=1}^{k}\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\prod_{l=1}^{j-1}\frac{M_{l}}{\pi_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\cdot$
$\quad\left(\underbrace{\mathbb{E}[\mu_{jm_{k}}^{k}\left(X,\overline{Z}_{j}\right)-\mu_{(j-1)m_{j}}\left(X,\bar{Z}_{j-1}\right)|X,\bar{Z}_{j-1},\overline{1}_{j}]}_{=0}\right)+\mu_{0}(X)\bigg{]}$
$=P\bigg{[}\mu_{0}(X)\bigg{]}.$
[…]
$=P\bigg{[}\sum_{j=0}^{k-1}\frac{A}{\hat{\pi}_{01}(X)}\left(\prod_{l=1}^{j-1}\frac{M_{j}}{\hat{\pi}_{l1}(X,\bar{Z}_{l})}\right)\cdot$
$\quad\big{(}\hat{\pi}_{j1}\big{(}X,\overline{Z}_{j}\big{)}-\pi_{j1}\big{(}X,\overline{Z}_{j}\big{)}\big{)}\big{(}\hat{\mu}_{jm_{k}}^{k}\big{(}X,\bar{Z}_{j}\big{)}-\mu_{jm_{k}}^{k}\big{(}X,\bar{Z}_{j}\big{)}\big{)}$
$\quad+\frac{A}{\hat{\pi}_{01}(X)}\left(\prod_{l=1}^{k-1}\frac{M_{j}}{\hat{\pi}_{l1}(X,\bar{Z}_{l})}\right)\cdot$
$\quad\big{(}\hat{\pi}_{km_{k}}\big{(}X,\overline{Z}_{j}\big{)}-\pi_{km_{k}}\big{(}X,\overline{Z}_{j}\big{)}\big{)}\big{(}\hat{\mu}_{km_{k}}^{k}\big{(}X,\bar{Z}_{k}\big{)}-\mu_{km_{k}}^{k}\big{(}X,\bar{Z}_{k}\big{)}\big{)}\bigg{]}$
[…]
where the last equality results from the positivity assumption that $\hat{\pi}_{k1}(X,\bar{Z}_{k})$ is bounded away from zero, for all $k\in[K]$, and from the Cauchy-Schwartz inequality. Then, assuming that the empirical process term is of order $o_{p}(n^{-1/2})$, we can write $\hat{\varphi}_{km_{k}}^{\text{EIF}}-\varphi_{km_{k}}$ as
$\hat{\psi}_{km_{k}}^{\text{EIF}}-\psi_{km_{k}}$ $=\mathbb{P}_{n}[m(O;\eta)-\psi_{km_{k}}]$
[…]
$\quad+o_{p}(n^{-1/2}).$
[…]
Turning next to to $\phi_{k}\triangleq\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]$, for all $k\in[K-1]$, we can similarly define $\gamma_{k}\left(X,\bar{Z}_{k}\right)$ iteratively as
$\gamma_{k}\left(X,\bar{Z}_{k}\right)$ $\triangleq\mathbb{E}\left[M_{k+1}\mid X,\bar{Z}_{k},\overline{1}_{k+1}\right]$
$\gamma_{j}\left(X,\bar{Z}_{j}\right)$ $\triangleq\mathbb{E}\left[\gamma_{j+1}\left(X,\bar{Z}_{j+1}\right)\mid X,\bar{Z}_{j},\overline{1}_{j+1}\right]\forall j\in[k-1].$
Similarly to the previous case, the EIF of $\phi_{k}$ is equal to
$\varphi_{k}(O)$ $=\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\prod_{l=1}^{k}\frac{M_{j}}{\pi_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\left(Y-\gamma_{k}\left(X,\bar{Z}_{k}\right)\right).$
$\quad+\sum_{j=0}^{k}\frac{A}{\pi_{01}\big{(}X\big{)}}\left(\prod_{l=1}^{j-1}\frac{M_{l}}{\pi_{l1}\big{(}X,\overline{Z}_{l}\big{)}}\right)\left(\gamma_{j}\left(X,\overline{Z}_{j}\right)-\gamma_{j-1}\left(X,\bar{Z}_{j-1}\right)\right)$
$\quad+\gamma_{0}(X)-\phi_{k}.$
Following similar arguments to the above, we have that
$\hat{\phi}_{k}^{\text{EIF}}-\phi_{k}$ $=\mathbb{P}_{n}[m_{2}(O;\eta)-\phi_{k}]$
[…]
$\quad+o_{p}(n^{-1/2}),$
where $m_{2}(O;\hat{\eta})=\varphi_{k}+\phi_{k}$. Thus, letting $R_{n}(k,\gamma)\triangleq\sum_{j=0}^{k}R_{n}\big{(}\hat{\pi}_{j})R_{n}\big{(}\hat{\gamma}_{j})$, $\hat{\phi}_{k}^{\text{EIF}}$ is consistent if $R_{n}(k,\gamma)=o(1)$ and it is semiparametric efficient if $R_{n}(k,\gamma)=o(n^{-1/2})$. This result implies that if all nuisance functions are consistently estimated and converge at faster than $n^{1/4}$ rates, then $\hat{\phi}_{k}^{\text{EIF}}$ is semiparametric efficient. I first establish the following lemma:
Lemma 1.
Let $X_{n}$ and $Y_{n}$ denote two convergent sequences, where $X_{n}=O_{p}(n^{-1/2})$ and $Y_{n}=o_{p}(n^{-1/2})$. Then, (a) $X_{n}Y_{n}=o_{p}(n^{-1/2})$, and (b) $X_{n}X_{n}=o_{p}(n^{-1/2})$.
Proof.
(a) $X_{n}=O_{p}(n^{-1/2})=n^{-1/2}O_{p}(1)=o_{p}(1)$. Thus, $X_{n}Y_{n}=o_{p}(1)o_{p}(n^{-1/2})=o_{p}(n^{-1/2})$. (b) $X_{n}X_{n}=O_{p}(n^{-1/2})O_{p}(n^{-1/2})=O_{p}(n^{-1})=n^{-1/2}O_{p}(n^{-1/2})=n^{-1/2}o_{p}(1)$ (by (a)). Thus, $X_{n}X_{n}=o_{p}(n^{-1/2})$. ∎
Using this lemma, I establish rate conditions for the semiparametric efficiency of $\eta_{k}=\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k}.$ We can analyze the asymptotic behavior of $\hat{\eta}_{k}=\hat{\tau}_{k-1}-\hat{\Delta}_{k-1}-\hat{\pi}_{k}\hat{\tau}_{k}$ via a distributional expansion of each plug-in estimator:
$\hat{\eta}_{k}$ $=\hat{\tau}_{k-1}-\hat{\Delta}_{k-1}-\hat{\pi}_{k}\hat{\tau}_{k}$
$=(\tau_{k-1}+\mathbb{P}_{n}[\tau_{k-1}^{\text{EIF}}]+\tau_{k-1}^{\text{EP}}+\tau_{k-1}^{\text{R2}}]-(\Delta_{k-1}+\mathbb{P}_{n}[\Delta_{k-1}^{\text{EIF}}]+\Delta_{k-1}^{\text{EP}}+\Delta_{k-1}^{\text{R2}}]$
$\quad-[\pi_{k}+\mathbb{P}_{n}[\pi_{k}^{\text{EIF}}]+\pi_{k}^{\text{EP}}+\pi_{k}^{\text{R2}}][\tau_{k}+\mathbb{P}_{n}[\tau_{k}^{\text{EIF}}]+\tau_{k}^{\text{EP}}+\tau_{k}^{\text{R2}}]$
$=(\tau_{k-1}+\mathbb{P}_{n}[\tau_{k-1}^{\text{EIF}}]+o_{p}(n^{-1/2})+\tau_{k-1}^{\text{R2}}]-(\Delta_{k-1}+\mathbb{P}_{n}[\Delta_{k-1}^{\text{EIF}}]+o_{p}(n^{-1/2})+\Delta_{k-1}^{\text{R2}}]$
$\quad-[\pi_{k}+\mathbb{P}_{n}[\pi_{k}^{\text{EIF}}]+o_{p}(n^{-1/2})+\pi_{k}^{\text{R2}}][\tau_{k}+\tau_{k}^{\text{EIF}}+o_{p}(n^{-1/2})+\tau_{k}^{\text{R2}}]$
$=[\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k}]+\mathbb{P}_{n}[(\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k})^{\text{EIF}}]$
$\quad+\tau_{k-1}^{\text{R2}}+\Delta_{k-1}^{\text{R2}}+\pi_{k}^{\text{R2}}+\tau_{k}^{\text{R2}}+O_{p}(n^{-1/2})O_{p}(n^{-1/2})+O_{p}(n^{-1/2})o_{p}(n^{-1/2})+o_{p}(n^{-1})+o_{p}(n^{-1/2})$
$=[\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k}]+\mathbb{P}_{n}[(\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k})^{\text{EIF}}]$
$\quad+\tau_{k-1}^{\text{R2}}+\Delta_{k-1}^{\text{R2}}+\pi_{k}^{\text{R2}}+\tau_{k}^{\text{R2}}+o_{p}(n^{-1/2}),$
where the penultimate equality follows from Proposition 1, and the final equality follows from Lemma 1.
Thus, for any $k\in\{1,\dots,K\}$, $\hat{\eta}_{k}=\hat{\tau}_{k-1}-\hat{\Delta}_{k-1}-\hat{\pi}_{k}\hat{\tau}_{k}$ is semiparametric efficient if $\tau_{k-1}^{\text{R2}}+\Delta_{k-1}^{\text{R2}}+\pi_{k}^{\text{R2}}+\tau_{k}^{\text{R2}}=o_{p}(n^{-1/2})$.
For the continuation terms ($\theta_{k}=(\Pi_{j=1}^{k}\pi_{j})\Delta_{k}+(\Pi_{j=1}^{k-1}\pi_{j})\eta_{k}$), I proceed by induction. Let $\hat{\Delta}_{*}=\Delta_{*}+\mathbb{P}_{n}[\Delta_{*}^{\text{EIF}}]+\Delta_{*}^{\text{EP}}+\Delta_{*}^{\text{R2}}$ and $\hat{\eta}_{*}=\eta_{*}+\mathbb{P}_{n}[\eta_{*}^{\text{EIF}}]+\eta_{*}^{\text{EP}}+\eta_{*}^{\text{R2}}$ be asymptotically linear, where $\in\{1,\dots K\}$. For $k=1$, we can asymptotically expand $\hat{\pi}_{1}\hat{\Delta}_{{}^{*}}$ as
$\hat{\pi}_{1}\hat{\Delta}_{*}$ $=(\pi_{1}+\mathbb{P}_{n}[\pi_{1}^{\mathrm{EIF}}]+\pi_{1}^{\mathrm{EP}}+\pi_{1}^{\mathrm{R}2})(\Delta_{*}+\mathbb{P}_{n}[\Delta_{*}^{\text{EIF}}]+\Delta_{*}^{\text{EP}}+\Delta_{*}^{\text{R2}})$
$=(\pi_{1}+\mathbb{P}_{n}[\pi_{1}^{\mathrm{EIF}}]+o_{p}(n^{-1/2})+\pi_{1}^{\mathrm{R}2})(\Delta_{*}+\mathbb{P}_{n}[\Delta_{*}^{\mathrm{EIF}}]+o_{p}(n^{-1/2})+\Delta_{*}^{\mathrm{R}2})$
[…]
and expand $(\Pi_{j=1}^{k-1}\pi_{j})\eta_{*}$, similarly, as
$\sum_{j=1}^{k-1}\hat{\pi}_{j}\hat{\eta}_{*}$ $=\eta_{*}+\mathbb{P}_{n}[\eta_{*}{}^{\text{EIF}}]+\tau_{*-1}^{\text{R2}}+\Delta_{*-1}^{\text{R2}}+\pi_{*}^{\text{R2}}+\tau_{*}^{\text{R2}}+o_{p}(n^{-1/2})$
$=\sum_{j=1}^{k-1}\pi_{j}\eta_{*}+\mathbb{P}_{n}[(\sum_{j=1}^{k-1}\pi_{j}\eta_{*})^{\text{EIF}}]$
$\quad+\tau_{*-1}^{\text{R2}}+\Delta_{*-1}^{\text{R2}}+\pi_{*}^{\text{R2}}+\tau_{*}^{\text{R2}}+\sum_{j\in\{1\dots,k^{*}-1\}:j\neq*}\pi_{j}^{\text{R2}}+o_{p}(n^{-1/2}),$
following a similar logic to above. Now, assume that, for $k^{*}\in\{1,\dots,K\}$, $(\Pi_{j=1}^{k^{*}}\hat{\pi}_{j})\hat{\Delta}_{*}=\Delta_{k}\prod_{j=1}^{k^{*}}\pi_{j}+\mathbb{P}_{n}[(\Delta_{*}\prod_{j=1}^{k^{*}}\pi_{j})^{\text{EIF}}]+\Delta_{*}^{\mathrm{R}2}+\sum_{j=1}^{k^{*}}\pi_{j}^{\mathrm{R}2}+o_{p}(n^{-1/2})$ and, further, that $(\Pi_{j=1}^{k^{*}-1}\hat{\pi}_{j})\hat{\eta}_{*}=\Pi_{j=1}^{k^{*}-1}\pi_{j}\eta_{*}+\mathbb{P}_{n}[(\Pi_{j=1}^{k^{*}-1}\pi_{j}\eta_{*})^{\text{EIF}}]+\tau_{*-1}^{\text{R2}}+\Delta_{*-1}^{\text{R2}}+\pi_{*}^{\text{R2}}+\tau_{*}^{\text{R2}}+\sum_{j\in\{1\dots,k^{*}-1\}:j\neq*}\pi_{j}^{\text{R2}}+o_{p}(n^{-1/2})$. Then, by induction, we have that
$(\Pi_{j=1}^{k^{*}+1}\hat{\pi})\hat{\Delta}_{*}$ $=\bigg{[}(\pi_{k^{*}+1}+\mathbb{P}_{n}[\pi_{k^{*}+1}^{\mathrm{EIF}}]+o_{p}(n^{-1/2})+\pi_{k^{*}+1}^{\mathrm{R}2})\bigg{]}$
$\quad\bigg{[}\Delta_{*}\prod_{j=1}^{k^{*}}\pi_{j}+\mathbb{P}_{n}[(\Delta_{*}\sum_{j=1}^{k^{*}}\pi_{j})^{\text{EIF}}]+\Delta_{*}^{\mathrm{R}2}+\sum_{j=1}^{k^{*}}\pi_{j}^{\mathrm{R}2}+o_{p}(n^{-1/2})\bigg{]}$
$=\Delta_{*}\prod_{j=1}^{k^{*}+1}\pi_{j}+\mathbb{P}_{n}[(\Delta_{*}\prod_{j=1}^{k^{*}+1}\pi_{j})^{\text{EIF}}]+\Delta_{*}^{\mathrm{R}2}$
$\quad+\sum_{j=1}^{k^{*}+1}\pi_{j}^{\mathrm{R}2}+O_{p}(n^{-1/2})O_{p}(n^{-1/2})$
$\quad+O_{p}(n^{-1/2})o_{p}(n^{-1/2})+o_{p}(n^{-1})+o_{p}(n^{-1/2})$
$=\Delta_{*}\sum_{j=1}^{k^{*}+1}\pi_{j}+\mathbb{P}_{n}[(\Delta_{*}\prod_{j=1}^{k^{*}+1}\pi_{j})^{\text{EIF}}]+\Delta_{*}^{\mathrm{R}2}+\sum_{j=1}^{k^{*}+1}\pi_{j}^{\mathrm{R}2}+o_{p}(n^{-1/2}),$
and that
$(\Pi_{j=1}^{k^{*}}\hat{\pi}_{j})\hat{\Delta}_{*}$ $=\bigg{[}(\pi_{k^{*}+1}+\mathbb{P}_{n}[\pi_{k^{*}+1}^{\mathrm{EIF}}]+o_{p}(n^{-1/2})+\pi_{k^{*}+1}^{\mathrm{R}2})\bigg{]}$
$\quad\bigg{[}\Pi_{j=1}^{k^{*}-1}\pi_{j}\eta_{*}+\mathbb{P}_{n}[(\Pi_{j=1}^{k^{*}-1}\pi_{j}\eta_{*})^{\text{EIF}}]+\tau_{*-1}^{\text{R2}}+\Delta_{*-1}^{\text{R2}}+\pi_{*}^{\text{R2}}+\tau_{*}^{\text{R2}}+\sum_{j\in\{1\dots,k^{*}-1\}:j\neq*}\pi_{j}+o_{p}(n^{-1/2})\bigg{]}$
$=\Pi_{j=1}^{k^{*}}\pi_{j}\eta_{*}+\mathbb{P}_{n}[(\Pi_{j=1}^{k^{*}}\pi_{j}\eta_{*})^{\text{EIF}}]+\tau_{*-1}^{\text{R2}}+\Delta_{*-1}^{\text{R2}}+\pi_{*}^{\text{R2}}+\tau_{*}^{\text{R2}}+\sum_{j\in\{1\dots,k^{*}\}:j\neq*}\pi_{j}^{\text{R2}}+o_{p}(n^{-1/2})$
It follows that, for any $k\in\{1,\dots,K\}$,
$(\Pi_{j=1}^{k}\hat{\pi}_{j})\hat{\Delta}_{k}=\Delta_{k}\sum_{j=1}^{k}\pi_{j}+\Pi_{j=1}^{k-1}\pi_{j}\eta_{k}+\mathbb{P}_{n}[(\Delta_{k}\sum_{j=1}^{k}\pi_{j}+\Pi_{j=1}^{k-1}\pi_{j}\eta_{k})^{\text{EIF}}]+\sum_{j=k-1}^{k}\Delta_{j}^{\mathrm{R}2}+\sum_{j=k-1}^{k}\tau_{j}^{\mathrm{R}2}+\sum_{j=1}^{k}\pi_{j}^{\mathrm{R}2}.$
Thus, $\hat{\theta}_{k}=(\Pi_{j=1}^{k}\hat{\pi}_{j})\hat{\Delta}_{k}+(\Pi_{j=1}^{k-1}\hat{\pi}_{j})\hat{\eta}_{k}$ is semiparametric efficient if $\sum_{j=k-1}^{k}\Delta_{j}^{\mathrm{R}2}+\sum_{j=k-1}^{k}\tau_{j}^{\mathrm{R}2}+\sum_{j=1}^{k}\pi_{j}^{\mathrm{R}2}=o(n^{-1/2})$. Proposition 2 then follows immediately by recognizing the rate conditions required for each of the constituent functionals of $(\pi_{k},\Delta_{k},\tau_{k})$ to be semiparametric efficient.
10 Derivation of RWR procedures
For simplicity, throughout the following I let $Z_{0}=X$ and $M_{0}=A$. I assume the following linear specification of the outcome model:
$\mathbb{E}[Y\mid\overline{Z}_{k},A,\overline{M}_{k}]$ $=\beta_{k,0}+c_{k,0}A+\sum_{j=1}^{k}\beta_{k,j}M_{j}+\eta_{k,1}^{\top}X^{\perp}+c_{k,1}AX^{\perp}+\sum_{j=1}^{k-1}\eta_{k,j}^{T}M_{j}X^{\perp}+\sum_{j=1}^{k}\gamma_{k,j}^{T}Z_{j}^{\perp}$ (17)
$\quad+\sum_{j=1}^{k-1}M_{j}\sum_{l=1}^{j}\xi_{k,k,l}^{\top}Z_{l}^{\perp},$
where $Z_{k}^{\perp}=Z_{k}-\mathbb{E}[Z_{k}|\overline{Z}_{k-1},M_{k-1}=1_{k-1}]$, $\forall k\in[0,\dots,K]$. In the following derivations, I use the fact that, $\forall k\in\{1,\dots,K\},$
$\int z_{k}^{\perp}dP(z_{k}|\overline{z}_{k-1},m_{k-1}=1)$
$=\mathbb{E}[Z_{k}-\mathbb{E}[Z_{k}|\overline{z}_{k-1},m_{k-1}=1]|\overline{z}_{k-1},m_{k-1}=1]$
$=0.$
Letting $X=Z_{0}$, the above also implies that $\int z_{0}^{\perp}dP(z_{0})=\mathbb{E}[Z_{0}-\mathbb{E}[Z_{0}]]=0$. Under sequential ignorability and assuming linearity of the outcome with respect to all antecedent variables, we have that
$\Delta_{k-1}$ $=\int\mathbb{E}[Y|\overline{M}_{k-1}=\overline{1}_{k-1},\overline{z}_{k},M_{k}=0]\prod_{j=0}^{k}dP(z_{j}|\overline{z}_{j-1},m_{j-1}=1)$
$-\int\mathbb{E}[Y|\overline{M}_{k-2}=\overline{1}_{k-2},\overline{z}_{k},M_{k-1}=0]\prod_{j=1}^{k}dP(z_{j}|\overline{z}_{j-1},m_{j-1}=1)$
$=\int\bigg{[}\beta_{k,0}+c_{k,0}+\sum_{j=1}^{k-1}\beta_{k,j}+\eta_{k,1}^{\top}X^{\perp}+c_{k,1}X^{\perp}+\sum_{j=1}^{k-2}\eta_{k,j}^{T}X^{\perp}+\sum_{j=1}^{k}\gamma_{k,j}^{T}Z_{j}^{\perp}+\sum_{j=1}^{k-2}\sum_{l=1}^{j}\xi_{k,k,l}^{\top}Z_{l}^{\perp})$
$-(\beta_{k,0}+c_{k,0}+\sum_{j=1}^{k-2}\beta_{k,j}M_{j}+\eta_{k,1}^{\top}X^{\perp}+c_{k,1}AX^{\perp}+\sum_{j=1}^{k-2}\eta_{k,j}^{T}X^{\perp}+\sum_{j=1}^{k}\gamma_{k,j}^{T}Z_{j}^{\perp}+\sum_{j=1}^{k-3}\sum_{l=1}^{j}\xi_{k,k,l}^{\top}Z_{l}^{\perp})\bigg{]}$
$\prod_{j=0}^{k}dP(z_{j}|\overline{z}_{j-1},m_{j-1}=1)$
$=\beta_{k,k-1}.$
Further, for $\tau_{k}\forall k\in\{1,\dots,K\}$ we have that
$\tau_{k}$ $=\int\mathbb{E}[Y|A=1,\overline{M}_{k}=\overline{1}_{k},\overline{z}_{k}]\prod_{j=0}^{k}dP(z_{j}|\overline{z}_{j-1},m_{j-1}=1)$
$-\int\mathbb{E}[Y|A=1,\overline{M}_{k-1}=\overline{1}_{k-1},\overline{z}_{k},M_{k}=0]\prod_{j=0}^{k}dP(z_{j}|\overline{z}_{j-1},m_{j-1}=1)$
$=\int\bigg{[}(\beta_{k,0}+c_{k,0}+\sum_{j=1}^{k}\beta_{k,j}+\eta_{k,1}^{\top}X^{\perp}+c_{k,1}X^{\perp}+\sum_{j=1}^{k-1}\eta_{k,j}^{T}X^{\perp}+\sum_{j=1}^{k}\gamma_{k,j}^{T}Z_{j}^{\perp}+\sum_{j=1}^{k-1}\sum_{l=1}^{j}\xi_{k,j,l}^{\top}Z_{l}^{\perp})$
$-(\beta_{k,0}+c_{k,0}+\sum_{j=1}^{k-1}\beta_{k,j}+\eta_{k,1}^{\top}X^{\perp}+c_{k,1}X^{\perp}+\sum_{j=1}^{k-2}\eta_{k,j}^{T}X^{\perp}+\sum_{j=1}^{k}\gamma_{k,j}^{T}Z_{j}^{\perp}+\sum_{j=1}^{k-2}\sum_{l=1}^{j}\xi_{k,j,l}^{\top}Z_{l}^{\perp})\bigg{]}$
$\prod_{j=0}^{k}dP(z_{j}|\overline{z}_{j-1},m_{j-1}=1)$
$=\beta_{k,k}.$
Finally, I assume that
$\begin{aligned} \mathbb{E}[M_{k+1}\mid A=1,\overline{Z}_{k},\overline{M}_{k}=\overline{1}_{k}]&=\theta_{0}+\sum_{k=0}^{k}\delta_{k+1}^{T}Z_{k}^{\perp}\end{aligned}.$
Then:
$\mathbb{E}[M(\overline{1}_{k+1})]$ $=\theta_{0}+\int\bigg{[}\sum_{k=0}^{k}\delta_{k+1}^{T}Z_{k}^{\perp}\prod_{j=0}^{k}dP(z_{j}|\overline{z}_{j-1},m_{j-1}=1)\bigg{]}$
$=\theta_{0}.$
11 Derivation of bias formulae for sensitivity analysis
In this section, I derive the bias formulae for the set ($\tau_{k},\Delta_{k}$) for all $k\in[K]$, where $K$ denotes the number of mediators considered in the decomposition, under a sequence of simplifying assumptions. Assume first that we have a binary unobserved confounder, $U$, for the treatment-outcome relationship. Assuming that $\alpha_{0}=\mathbb{E}[Y|x,a,U=1]-\mathbb{E}[Y|x,a,U=0]$ does not depend on $x$ or $a$, and further that $\beta_{0}=\text{Pr}[U=1|x,A=1]-\text{Pr}[U=1|x,A=0]$ does depend on $x$, for $\tau_{0}=\mathbb{E}[Y(1)-Y(0)]\triangleq\text{ATE}$, I then have that $\text{bias}(\tau_{0})=\alpha\beta$ [VanderWeele and Arah, 2011].
Next, consider an unobserved confounder, $U_{k}$ that affects both $M_{k}$ and $Y$ for any $k\in\{1,\dots,K\}$. Then, under a weaker iteration of Assumption 4 (Sequential Ignorability), i.e.,
$Y(\overline{1}_{k},m_{k})\perp\!\!\!\perp(A,\overline{M}_{k})|X,A,U_{k},\overline{Z}_{k},\overline{M}_{k-1}\forall k\in[K],$$\mathbb{E}[Y(\overline{1}_{k},m_{k})]$ is identified as
$\mathbb{E}[Y(\overline{1}_{k},m_{k})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},u_{k}]\big{[}dP(u_{k}|x,\overline{z}_{k},\overline{1}_{k-1})\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x).$
By contrast, under Assumption 4, my estimator of $\mathbb{E}[Y(\overline{1}_{k},m_{k})]$, $\tilde{\mathbb{E}}[Y(\overline{1}_{k},m_{k})]$, converges to
$\tilde{\mathbb{E}}[Y(\overline{1}_{k},m_{k})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},u_{k}]\big{[}dP(u_{k}|x,\overline{z}_{k},\overline{1}_{k})\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x).$
I invoke the following three assumptions: (Assumption $A_{k}$) $\alpha_{k}=\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=1]-\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=0]$ does not depend on ($x,\overline{z}_{k},\overline{1}_{k},m_{k})$; (Assumption $B_{k}$) $\beta_{k}=\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},m_{k}]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}$ does not depend on ($x,\overline{z}_{k}$); Assumption ($C$) $U_{k}$ is binary. Taking the difference between the quantities in the above two equations thus gives that, for any $m_{k}\in\{0,1\}$, for any $k\in[K]$, we have that
$\text{bias}(\tilde{\mathbb{E}}[Y(\overline{1}_{k},m_{k})])$ $=\int\big{(}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=1]-\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k},U_{k}=0]\big{)}\cdot$
$\quad\big{(}\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},m_{k}]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x).$ (18)
Consider first $\text{bias}(\Delta_{k-1})=\text{bias}(\tilde{\mathbb{E}}[Y(\overline{1}_{k},0)-Y(\overline{1}_{k-1},0)])$. Under mediator monotonicity (Assumption 1), I immediately have that $\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},m_{k}]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}$, and thus that $\text{bias}(\Delta_{k-1})=\text{bias}(\tilde{\mathbb{E}}[Y(\overline{1}_{k},0)])$, which can be written as
$\text{bias}(\tilde{\mathbb{E}}[Y(\overline{1}_{k},0)])$ $=\int\big{(}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},0,U_{k}=1]-\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},0,U_{k}=0]\big{)}$
$\quad\big{(}\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},0]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x)$
$=\int\big{(}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},0,U_{k}=1]-\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},0,U_{k}=0]\big{)}\cdot$
$\quad\big{(}\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},0]-\big{(}\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k+1}]\text{Pr}[M_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]$
$+\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},0]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},0]\text{Pr}[M_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}\big{)}$
$\quad\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x)$
$=-\int\big{(}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},0U_{k}=1]-\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},0_{k},U_{k}=0]\big{)}\cdot$
$\quad\big{(}\big{(}\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k+1}]-\text{Pr}[U_{k}=1|x,\overline{z}_{k},\overline{1}_{k},0]\big{)}\text{Pr}[M_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\big{)}$
$\quad\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x).$
Next, applying assumptions $A_{k}$ and $B_{k}$, we can write
$\text{bias}(\tilde{\mathbb{E}}[Y(\overline{1}_{k+1},0)])=-\alpha_{k}\beta_{k}\int_{x}\int_{\overline{z}_{k}}\text{Pr}[M_{k}=1|x,\overline{z}_{k},\overline{1}_{k}]\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x).$
Second, to compute $\text{bias}(\tau_{k})=\text{bias}(\mathbb{E}[Y(\overline{1}_{k+1})-Y(\overline{1}_{k},0)])$ for any $k\in\{1,\dots K\}$, beginning with Equation 18 and applying assumptions $A_{k}$ and $B_{k}$ once again, we have that
$\text{bias}(\tau_{k})$ $=\alpha_{k}\beta_{k}.$